### v1

# Robust Denoising Neural Reranker for Recommender Systems

Abstract
For multi-stage recommenders in industry, a user request would first trigger a simple and efficient retriever module that selects and ranks a list of relevant items, then calls a slower but more sophisticated deep reranking model that refines the item arrangement before exposure to the user. The latter model typically reranks the item list conditioned on the users’ history content and the initial ranking from retrievers. Although this two-stage retrieval-ranking framework demonstrates practical effectiveness, the significance of retriever scores from the previous stage has been limitedly explored, which is informative. In this work, we first theoretically analyze the limitations of using retriever scores as the reranker’s input directly and argue that the reranking task is essentially a noise reduction problem from the retriever scores. Following this notion, we derive an adversarial framework, DNR, that associates the denoising reranker with a carefully designed noise generation module. We extend the conventional score error minimization term with three augmented objectives, including: 1) a denoising objective that aims to denoise the noisy retriever scores to align with the user feedback; 2) an adversarial retriever score generation objective that improves the exploration in the retriever score space; and 3) a distribution regularization term that aims to align the distribution of generated noisy retriever scores with the real ones. Extensive experiments are conducted on three public datasets, together with analytical support, validating the effectiveness of the proposed DNR.
1 Introduction
Recommender systems aim to personalize a list of items that maximizes user engagement signals such as purchases, video plays, and link sharing. To achieve this, industrial solutions [1, 2, 3] typically employ a multi-stage architecture comprising a retriever and a reranker. The retriever, which may further decompose into several consecutive sub-stages of retrieval and ranking models, efficiently filters a large item pool down to a manageable candidate set of relevant items. Then the reranker adopts a more sophisticated but more accurate model that refines the candidate item set into an optimal ordering for the list-wise exposure. Formally, the goal of the latter reranking task [4, 5, 6, 7] is to learn an ordering function over the retrieved candidates that maximizes user responses on the top-reranked items. This task is particularly challenging due to the combinatorial nature of the item permutation space. Exhaustively searching for the optimal list exposure in this space becomes infeasible as the candidate set grows, making this a key bottleneck in real-world recommender systems.
To overcome this challenge, early studies propose to refine the initial retriever scores [5, 8, 6] by modeling the item mutual influences in the candidate item list. Recent state-of-the-art solutions find it more effective to frame the reranking task as a generative modeling task [9, 10, 11] where the list generation probability should directly align with the user’s list-wise rewards. Yet, these methods largely overlook the potential of initial retriever scores, which may offer rich prior information to enhance the reranking performance. In practice, a straightforward approach to leverage initial retriever scores is to include them as input features directly, which shows promise to improve reranking performance as in Figure 1(a).
Reranking as a noise reduction problem: Though effective, this naive solution neglects the noisy nature of the initial retriever scores, which may cause model robustness concerns and downgrade the effectiveness of the learned reranking model. Specifically, the retriever stage, constrained by strict latency and computational budgets, usually adopts a simpler model than the reranker. In contrast, the reranker, operating on a much smaller candidate set, can use a more sophisticated architecture and thus achieves significantly higher accuracy (Figure 1(b)). In other words, by assuming higher noise in the retriever scores, it naturally indicates a noise reduction process of the reranking stage. This noise reduction pattern is more evident when we observe the changes in the embedding representations of ground truth target items, comparing Figure 1(c) and Figure 1(d). This means that the reranker behaves like a denoising model even without explicit modeling of the retriever’s noise. However, as we will show in section 3.1, the retriever’s noise theoretically contributes to the misalignment between the traditional objective of reranking and the user response likelihood maximization goal, restricting the generalization ability of the reranker. As a countermeasure, we need a more robust solution that can learn a denoising reranker that is aware of the retriever’s noisy score distribution, while taking advantage of the information from retriever scores.
(a) Initial retriever scores further boost the performance.
(b) Performance comparison of two stages in MovieLens dataset.
(c) Target item and selected candidates embedding in the retriever stage.
(d) Noise reduction process in the item embedding perspective.
Figure 1: (a) The informative retriever score and (b-d) noise reduction nature of reranking. The reranker in this example uses transformer model to select top-20 items from the candidates. “Rerank w/ s” represents using retriever scores as additional item features for reranker. The blue shaded circles in (c) and (d) represent the retriever’s selection of candidate items. The green shaded circle in (d) represents the reranker’s selection of the items of exposure.
The proposed solution: Based on the previous analysis, we propose to augment the denoising neural reranker (DNR) with a carefully designed noise injection model, which estimates the initial retriever score’s posterior distribution. We theoretically show that the data likelihood maximization goal derives a solution framework that includes three learning objectives: 1) an augmented denoise objective that aims to boost the user feedback reconstruction under unseen retriever scores, increasing the reranker’s effectiveness to denoise noisy retriever scores; 2) an adversarial retriever score generation objective that aims to generate irregular retriever scores for the reranker that are hard to denoise, which in turn improves the reranker’s robustness; and 3) a score posterior regularization term that aims to align the noise-aware posterior model with the initial retriever score distribution, guaranteeing data alignment with the system behavior in previous stages. Additionally, we show that the noise-injecting posterior and the denoising reranker collaborate with each other in a similar way to the one-step diffusion framework [12], but standard Gaussian noise is not necessarily an empirically optimal choice. Instead, we illustrate that the binomial nature of the user response better suits a Beta distribution in the retriever score space, which derives our heuristic solution for the posterior. Still, this naive design lacks personalization abilities. To mitigate this limitation, we design a learnable posterior model that approximately optimizes the three augmented learning objectives.
Empirical supports and findings: We conduct empirical experiments on three public datasets to verify the superiority of DNR against leading rerankers. Notably, the reranker paired with the learnable posterior best aligns with the observed ground-truth posterior, which theoretically improves the alignment between the reranker’s learning objective and the user feedback data likelihood maximization goal, resulting in the best empirical reranking performance in our experiment.
In general, this work provides evidence that the initial retriever scores may be more informative than we think, and a more comprehensive utilization of this information may derive a solution that surpasses state-of-the-art methods.
2 Related Work
2.1 Multi-stage Recommendation
Modern content-intensive web services with large candidate pools typically employ multi-stage recommender systems [13, 14, 7] (also known as cascade ranking [15, 16] process) to balance accuracy and efficiency under strict latency constraints. Each stage progressively narrows down the candidate set while employing increasingly sophisticated models. While most existing work focuses on independently optimizing individual stages (e.g., retrieval [17, 18] or reranking stage [7, 6]), recent studies explore joint optimization across the entire pipeline [16, 13, 19]. These approaches leverage later-stage models to guide earlier-stage learning, improving overall coherence and performance [20, 21]. However, our work takes a complementary direction—–using earlier-stage scores to regulate later-stage optimization.
2.2 Reranking Strategies
In sequential recommendation, re-ranking [6, 22, 8] aims to refine the initial ranking list generated by a previous retrieval stage. Existing re-ranking approaches can be broadly categorized into the following paradigms. Single-point re-scoring methods [23, 24] independently predict a refined score for each item and re-rank them accordingly. While efficient, they ignore dependencies between items, potentially leading to suboptimal list-level performance. Unlike single-point approaches, list-refinement methods [6, 8, 25] explicitly model mutual influences between items by treating the pre-ranked candidate list as input. Techniques such as Transformer-based architectures [6] capture user preference to optimize the list holistically. Generator-evaluator methods [22, 9] first generate multiple candidate lists and then evaluate them using a learned utility function to select the best one. They can effective generate a high-quality recommendation list but suffer from the increasing computational cost due to the generation-evaluation loop. Emerging techniques leverage diffusion models [11, 26] to generate refined item lists by iteratively denoising a ranking distribution, which excel in capturing complex user preferences.
2.3 Adversarial Learning
Adversarial learning [27] has emerged as a powerful technique for data augmentation in recommendation systems, improving the models robustness and generalization. By leveraging generative adversarial networks (GANs) [18] or adversarial training [28, 29], these methods synthesize high-quality user-item interactions or perturb existing data to enhance model robustness. For instance, IRGAN [18] employs a minimax game between a generator (sampling hard negative items) and a discriminator (distinguishing real vs. synthetic interactions) to improve ranking performance. Similarly, AdvIR [30] introduces adversarial perturbations to embedding spaces, forcing the model to learn more generalizable representations. These methods demonstrate that adversarial learning can not only augment training data but also enhance models’ robustness. Our method is designed to add noise by adversarial learning, which can enhance the denoise models (rerankers)’ robustness to uncertain noise in the pre-ranking scores.
3 Methodology
Figure 2: Overall framework of multi-stage recommender system (on the left) and the noise reduction formulation of our method, DNR.
3.1 Problem Formulation
For a given user request $\mathbf{u}$ (which may contain information about the candidate items $\mathcal{I}(u)$, user profile features, user interaction history), we observe the following during data collection: The candidate set $\mathcal{I}(u)=\{i_{1},i_{2},\cdots,i_{n}\}$ retrieved by the retriever; The corresponding (continuous) retriever scores of $\mathcal{I}(u)$ are $\mathbf{x}_{u}=[x_{i_{1}},x_{i_{2}},\cdots,x_{i_{n}}]\in[0,1]^{n}$, which essentially describes the use feedback likelihood in the view of the retriever; The ground truth user feedback (i.e., binary labels such as click, watch, share, etc.) of the candidate set, i.e., $\mathbf{z}_{u}=[z_{i_{1}},z_{i_{2}},\cdots,z_{i_{n}}]\in\{0,1\}^{n}$. Note that for recommender systems that further select K items after the reranking process, we may not observe the user feedback for the remaining items. Yet, from a probabilistic viewpoint, we can assume zeros as defaults since the item passing the reranking process is the premise of observing the user feedback signals.
The Noise Reduction Task: In our formulation, we denote the initial retriever as the prior $p_{\text{x}}(\mathbf{x}|\mathbf{u})$ and the learnable reranker $q_{\theta}(\mathbf{z}|\mathbf{x},\mathbf{u})$ as the conditional user feedback likelihood estimator . The $q_{\theta}$ predicts the user feedback given the retriever scores and the user context, i.e., $\hat{\mathbf{z}}_{u}\sim q_{\theta}(\mathbf{z}|\mathbf{x}_{u},\mathbf{u})$, and the learning goal adopts the general data likelihood maximization objective:
$\max p(\mathbf{z}_{u}|\mathbf{u})$ (1)
where the ground-truth $p(\mathbf{z}_{u}|\mathbf{u})$ is determined by the retriever $p_{\text{x}}$, the reranker $q_{\theta}$, and the user.
Note: For simplicity of the presentation, we omit the user condition term $\mathbf{u}$ for the remainder of the paper (e.g. we will denote the reranker as $q_{\theta}(\mathbf{z}|\mathbf{x}_{u})$ and simplify the learning goal as $\max p(\mathbf{z}_{u})$), since we adopt an analysis framework on the user request level.
3.2 Analytical Limitation of Direct Optimization Methods
As we have mentioned in section 1, involving retrieval scores as input is beneficial for existing reranking methods. Regardless of the reranker design, the corresponding objective function is:
$\mathcal{L}_{\text{direct}}=-\mathbb{E}_{\mathbf{x}\sim p_{\text{x}}}[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})]$ (2)
which maximizes the user behavior alignment conditioned on the given retrieval scores. In practice, according to the binary nature of $\mathbf{z}_{u}$, we can simplify the learning objective as standard binary cross-entropy (BCE) loss. Theoretically, we can find that $\mathcal{L}_{\text{direct}}$ optimizes an upper bound of the negative data log likelihood (details in Appendix A.2):
$-\log p(\mathbf{z}_{u})$ $=\mathcal{L}_{\text{direct}}+L_{1}+L_{2}$ (3)
$L_{1}$ $=\mathbb{E}_{\mathbf{x}\sim p_{\text{x}}}\Big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})}\Big]$
$L_{2}$ $=-D_{\text{KL}}\big(p_{\text{x}}(\mathbf{x})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big).$
where $p_{\text{z|x}}$ represents the ground truth user feedback probability conditioned on retriever scores, and $p_{\text{x|z}}$ represents the ground truth posterior. Intuitively, smaller values of $L_{1}$ and $L_{2}$ make the solution of Eq.(2) more aligned with the goal of Eq.(1). However, these two terms are not fully controlled by the direct optimization framework and may generate a non-robust solution of $q_{\theta}$:
Irregular cases from retriever: The expectation term $L_{1}$ describes the difference between $q_{\theta}$ and the ground truth $p_{\text{z|x}}$ under the retriever prior. This term is non-positive and is minimized when $q_{\theta}(\mathbf{z}|\mathbf{x})=p(\mathbf{z}|\mathbf{x})$ for any variables. Though $L_{1}$ is expected to be small when the reranker’s predictions align with the user feedback in dominant cases, the retriever may happen to generate an $\mathbf{x}_{u}$ that the reranker fails to comprehend, inducing a significant divergence between the optimization of Eq.(2) and the learning goal of Eq.(1).
Regular cases with noisy retriever scores: The KL term $L_{2}$ describes the difference between the retriever score distribution $p_{\text{x}}$ and the obtained posterior given the knowledge of $\mathbf{z}_{u}$. This term is also non-positive and can be non-trivial due to the noisy nature of the retriever (as illustrated in Figure 1). This means that there might exist consistently significant misalignment between Eq.(2) and Eq.(1).
3.3 Implementing Noise Injection Posterior
To overcome the aforementioned limitations, we propose to complement the denoising reranker $q_{\theta}$ with a generative posterior $p_{\phi}(\mathbf{x}|\mathbf{z}_{u})$ that explicitly models the ground truth retriever score posterior distribution $p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})$. As described in Section 1, $p_{\text{x|z}}$ behaves as a noise-injecting model, so:
$\mathbf{x}_{u}^{\prime}=(1-\lambda_{c})\mathbf{z}_{u}+\lambda_{c}\bm{\epsilon}_{\phi},$ (4)
where the sampled error term $\bm{\epsilon}_{\phi}$ represents the noisy behavior of the posterior. To this end, we investigate two choices of noise samplers for $\bm{\epsilon}_{\phi}$:
  • Heuristic Sampler: This option uses a non-parametric sampler $\bm{\epsilon}_{\phi}^{\text{heuristic}}$ with a manually designed distribution. For example, we can adopt standard Gaussian noises similar to [31, 26], but we notice that this is not a good representation of the ground truth. Intuitively, given the binary user response signals, a natural assumption of the retriever prior is the Beta distribution, where the parameters depend on the user request $\mathbf{u}$. What follows is a Beta posterior distribution, which derives our heuristic solution as $\bm{\epsilon}_{\phi}\sim\text{Beta}(\alpha,\beta)$. For simplicity, we set $\alpha=\beta=0.5$ to approximate the uncertainty, which would help improve the chance of generating harder samples for the reranker and consequently improve the robustness of the solution. Note that this implementation assumes that the noise does not depend on the user, and the personalization only comes from the label $\mathbf{z}_{u}$;
  • Model-based Sampler: While the heuristic sampler is effective in our empirical study, we are skeptical on whether the fixed setting of $\alpha$ and $\beta$ is optimal for all user requests. To achieve a personalized noise estimation, we use a learnable model $\bm{\epsilon}_{\phi}^{\text{model}}$ with parameter $\phi$ that will be optimized along with the reranker (details in the next section). As we will show in Section 4.3, this choice is empirically better, and the learned error exhibits a Beta distribution without explicitly defining it. Without specification, our standard DNR solution uses this model-based sampler.
3.4 Overall Learning Framework
In this section, we show how to optimize the user feedback alignment i.e., Eq.(1) by collaborating the learning of the reranker and the discovery of better noise-injecting posterior. Specifically, following a variational inference framework, we can decompose the negative log-likelihood objective into three loss terms (detailed derivation in Appendix A.1):
$-\log p(\mathbf{z}_{u})$ $=\mathcal{L}_{\text{z}}+\mathcal{L}_{\text{x}}+\mathcal{L}_{\text{adv}}+\delta_{\text{x}}$ (5)
$\mathcal{L}_{\text{z}}$ $=-\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\Big[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})\Big]$
$\mathcal{L}_{\text{adv}}$ $=\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\Big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})}\Big]$
$\mathcal{L}_{\text{x}}$ $=D_{\text{KL}}\Big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\Big),$
which consists of an augmented feedback reconstruction loss term $\mathcal{L}_{\text{z}}$, an adversarial noise exploration loss $\mathcal{L}_{\text{adv}}$, and a posterior distribution regularization term $\mathcal{L}_{\text{x}}$. The additional term $\delta_{\text{x}}=-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)\leq 0$ represents the divergence between the noise injection posterior $p_{\phi}$ and the corresponding ground truth $p_{\text{x|z}}$. Intuitively, we expect to achieve a small divergence in $\delta_{\text{x}}$ and $L_{\text{x}}$ so that the optimization of the three terms is aligned with the goal of Eq.(1).
Denoising with Augmented Retriever Scores: $\mathcal{L}_{\text{z}}$ in Eq.(5) denotes the reconstruction loss under the noisy retriever scores sampled by the posterior. Specifically, we can first sample $\mathbf{x}_{u}^{\prime}\sim p_{\phi}(\cdot|\mathbf{z}_{u})$ as the generated retriever scores, then pass them into the reranker $q_{\theta}$ for the reconstruction of $\mathbf{z}_{u}$. The optimization can also be simplified into the BCE loss with $\mathbf{z}_{u}$ as labels, and it only updates $\theta$. Note that the generated scores $\mathbf{x}_{u}^{\prime}$, different from the observed $\mathbf{x}_{u}$, essentially serve as the augmented scores that the initial retriever would have produced given the user request. This means that the reranker will have to learn a more robust reconstruction model of the ground-truth user feedback. Combining with the direct optimization, we have the final reconstruction loss for the reranker:
$\mathcal{L}_{\theta}=\mathcal{L}_{\text{direct}}+\lambda_{m}\mathcal{L}_{\text{z}},$ (6)
where $\lambda_{m}$ controls the magnitude of the augmentation, and the posterior $p_{\phi}$ is fixed.
Adversarial Noise Learning: $\mathcal{L}_{\text{adv}}$ describes how the reranker behaves differently from the true user feedback under the viewpoint of the noise-injecting posterior $p_{\phi}$. In practice, this term is expected to be small since the ratio $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})\approx 1$ in major cases due to the user feedback alignment of Eq.(6) and the augmented objective $\mathcal{L}_{\text{z}}$. Nevertheless, in the view of the posterior, we can see that $\mathcal{L}_{\text{adv}}$ encourages adversarial samples of $\mathbf{x}$ that produce small ratios of $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$. In other words, a better $p_{\phi}$ should be able to find irregular samples that are correctly recovered by the true distribution $p_{\text{z|x}}$ but are incorrectly reconstructed by the reranker $q_{\theta}$. For the posterior with heuristic sampler $\bm{\epsilon}_{\phi}^{\text{heuristic}}$, it is not optimizable once $\alpha$ and $\beta$ are chosen. In contrast, for the model-based sampler $\bm{\epsilon}_{\phi}^{\text{model}}$, we follow the intuition of irregular case exploration, and approximate $\mathcal{L}_{\text{adv}}$ by first fixing the reranker and then optimizing the posterior with a reconstruction minimization loss:
$\min_{\phi}\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u}^{\prime}).$ (7)
In our design, we directly output $\mathbf{x}_{u}^{\prime}$ using $\phi$ so that the end-to-end learning becomes feasible for Eq.(7). In cases where the sampling process is not directly optimizable, one may have to consider other alternatives like reparameterization tricks [32] or policy gradient [33].
On the other hand, $\mathcal{L}_{\text{adv}}$ does not encourage over-exploration of $\mathbf{x}_{u}^{\prime}$, which may negatively impact the learning of the reranker. In such cases, the generated scores are unrealistically different from the true posterior, diverging from the user behavior, inducing small ground truth $p_{\text{z|x}}$ and large ratio of $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$, violating $\mathcal{L}_{\text{adv}}$.
Noise Regularization: $\mathcal{L}_{\text{x}}$ estimates the distributional divergence between the posterior estimator $p_{\phi}$ and the prior $p_{\text{x}}$, and we aim to find a $p_{\phi}$ that reduces this divergence. Intuitively, this term favors the posteriors that can mimic the noisy behavior of the prior $p_{\text{x}}$ and remove the additional accuracy-related information from $\mathbf{z}$. We can calculate this term by first sampling retriever scores as $\mathbf{x}^{\prime}_{u}\sim p_{\phi}(\cdot|\mathbf{z}_{u})$, then estimating their log ratio difference from that of the prior. In cases where we do not have the model to directly estimate $p_{\text{x}}(\mathbf{x}_{u}^{\prime})$, one can circumvent this issue by directly aligning $\mathbf{x}_{u}^{\prime}$ with the observed samples $\mathbf{x}_{u}$:
$\mathcal{L}_{\text{x}}^{\prime}\approx-\mathbb{E}_{p_{\phi}}\big[\log\frac{\mathbf{x}_{u}}{\mathbf{x}^{\prime}_{u}}\big],$ (8)
which is an instance-level objective more rigorous than $\mathcal{L}_{\text{x}}$. Intuitively, a smaller value of $\mathcal{L}_{\text{x}}^{\prime}$ means that $\phi$’s output is close to the prior’s output $\mathbf{x}_{u}$, which restricts the distribution-level objective $\mathcal{L}_{\text{x}}$ to small values. Yet, we remind readers that it is not the case vice versa. Additionally, it is infeasible to repeat the exact same request and obtain augmented samples to cover the prior distribution of $p_{\text{x}}$. Fortunately, the collaborative learning across different user requests will approximate the sampling process of pair $(\mathbf{x}^{\prime}_{u},\mathbf{x}_{u})$ with similar patterns.
Again, note that the heuristic sampler $\bm{\epsilon}_{\phi}^{\text{heuristic}}$ is not optimizable by this objective, but one can achieve a small value of $\mathcal{L}_{\text{x}}$ by carefully designing the distribution (e.g., choosing Beta instead of Gaussian). For the training of the posterior with model-based sampler $\bm{\epsilon}_{\phi}^{\text{model}}$, we first adopt the heuristic sampler in the first $\lambda_{e}$ epochs to ensure learning stability, then engage the adversarial learning of $\mathcal{L}_{\text{adv}}$ and the optimization of Eq.(8) afterwards, switching the sampler to $\bm{\epsilon}_{\phi}^{\text{model}}$. As the summary of our solution, we present our training procedure in Appendix B.
4 Experiments
In this section, we conduct extensive experiments to answer the following questions:
RQ1: How does DNR perform in the reranking tasks of recommendation systems compared with leading baselines? RQ2: How well is the denoising formulation compared to alternative methods of leveraging retriever scores? RQ3: What is the performance difference between Gaussian noise and Beta noise? RQ4: How does the model-based sampler compare to the heuristic sampler in practice? RQ5: How does different objectives in DNR contribute to the performance? RQ6: How sensitive is DNR to the hyperparameters $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$?
4.1 Expermental Settings
Datasets. We conduct experiments on three real-world datasets, including ML-1M [34], Kuaivideo [3], and Amazon-books [35]. The detailed description and the statistics of the processed datasets for both the retriever and the reranker stages are presented in Appendix C.1.
Baselines: We compare the performance of DNR with four categories of recommendation methods in our reranking task, including traditional recommenders (i.e., SASRec [23], Caser [36], GRU4Rec [37], and MiDNN [24]), which individually predict the scores of candidate items and rank them accordingly; list-refinement methods (i.e., DLCM [8], SetRank [38], PRM [6], and MIR [25]), which encode the mutual inference in candidate items and refine the rankings based on this information; generator-evaluator methods (i.e., EGRerank [22], Pier [9], and NAR4Rec [10]), which generate multiple reranked item lists and evaluate them to select the best one for users; and diffusion-Based methods (i.e., DiffuRec [26] and DCDR [11]) which utilize diffusion models to generate items or lists for recommendation. Detailed information for our baselines is presented in Appendix C.2.
Implementation Details: Our experiments are implemented on Nvidia A40 GPUs with Python 3.8 and PyTorch 1.12. We construct the retriever stage with the collaborative filtering method [39] to obtain the initial retriever scores for the top-50 candidate items. The detailed setting and experimental results of the retriever stage are presented in Appendix D.1. The reranker reorders these items filtered by retrievers and selects the top $K=6$ for display. The dimension of rerankers’ hidden embedding is 128 across all models, and all models are trained until convergence. The learning rate and weight decay are tuned within the range of $[0.01,0.001,0.0001]$ and $[0,0.1,0.01,0.001]$, respectively. We tune the hyperparameters with $\lambda_{c}\in[0.1,1.0]$, $\lambda_{m}\in[0.1,1.0]$, and $\lambda_{e}\in[0,200]$.
Evaluation Metrics: Following common practices, we employ Hit Ratio@6 (H@6), NDCG@6 (N@6), MAP@6 (M@6), F1@6, and AUC for performance evaluation of the reranker stage. Although not the focus of this paper, we adopt the AUC metric to evaluate the performance of the retriever stage. And the retriever is fixed during training of rerankers.
4.2 Main Results (RQ1)
We compare the performance of our proposed DNR with multiple leading baselines and the results are detailed in Table 1. “DNR-G” and “DNR-B” denote two alternatives to implement $\bm{\epsilon}_{\phi}^{\text{heuristic}}$ by injecting Gaussian noise and Beta noise, respectively, before switching to the model-based sampler. We can observe that the generator-evaluator-based methods (i.e., EGRerank [22], Pier [9], and NAR4Rec [10]) achieve comparable performance among the multiple baselines, validating the effectiveness of the generator-evaluator paradigm in the reranker stage. In comparison, our DNR solutions consistently outperform the state-of-the-art baselines across the three benchmarks. This highlights the effectiveness of our noise reduction formulation and the significance of improving reranker robustness in our solution.
Table 1: Overall performance of different methods for the reranking recommendation. The best scores are in bold and the best baseline scores are underlined, respectively. All improvements are statistically significant with student t-test $p<0.05$.
Methods ML-1M Kuaivideo Book
H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC
SASRec 59.79 72.16 61.84 65.41 85.98 36.72 54.01 41.66 43.00 77.86 60.27 69.32 58.17 62.44 83.64
Caser 58.60 71.34 60.55 64.14 86.53 36.66 54.04 41.65 43.93 78.39 59.92 69.31 58.06 62.09 84.21
GRU4Rec 58.07 69.91 59.03 63.47 86.74 37.28 54.39 42.20 43.66 74.67 53.69 59.24 46.71 55.65 81.19
MiDNN 56.86 70.30 59.28 62.16 86.87 37.32 54.56 42.38 43.72 74.73 60.28 69.61 58.58 62.45 83.02
DLCM 62.31 73.87 63.82 67.96 89.35 39.69 60.67 48.90 46.61 75.80 66.80 75.88 65.39 69.28 91.93
SetRank 59.35 73.10 62.51 64.72 88.84 44.75 64.39 51.88 52.59 89.93 66.15 75.70 64.84 68.64 92.01
PRM 60.09 72.85 62.21 65.51 88.20 39.92 55.93 42.97 46.18 85.15 67.86 76.88 66.44 70.42 92.00
MIR 62.22 74.33 64.47 67.97 87.76 37.01 55.79 43.16 44.50 79.95 66.08 71.48 56.62 68.62 91.82
EGRank 62.76 74.75 64.97 68.46 88.72 40.09 59.01 47.52 47.06 77.44 70.73 80.75 72.40 73.33 89.40
Pier 62.74 75.99 65.98 68.74 90.43 45.35 65.11 52.55 53.35 90.93 71.14 80.22 71.62 73.74 92.26
NAR4Rec 62.81 75.01 65.42 68.31 88.30 44.31 63.83 51.45 52.08 89.94 70.08 79.46 70.69 72.66 92.44
DiffuRec 62.16 75.80 66.28 67.83 91.48 40.78 61.16 50.77 48.00 81.08 58.62 68.77 57.25 60.79 84.74
DCDR 60.79 73.21 62.57 66.58 89.01 43.86 63.52 50.92 51.66 89.92 65.84 70.36 54.76 67.33 88.07
[…]
4.3 Ablation Study (RQ2, RQ3, RQ4, & RQ5)
RQ2: To investigate different ways to leverage retriever scores, we develop three PRM variants for the noise generator: “c score” refers to leveraging a concatenate operation to combine the score features with encoded user states from users’ interaction history; “+ score” denotes incorporating the score feature with user states by a simple addition; “w score” represents adopting the initial retriever scores as the weights of candidate items. The frameworks of the three variants are shown in Appendix C.3. We report the experimental results in Table 2. The variants “c score”, “+ score”, and “w score” achieve better performance than the PRM backbone and slightly better performance than the state-of-the-art baseline Pier, validating that it is beneficial to include the information from retriever scores in the reranking stage. Additionally, our DNR solutions outperform all three variants, indicating a superior strategy to leverage the retriever scores.
Table 2: Ablation Study for the different ways of leveraging the retriever scores. Performance that exceeds Pier’s is indicated with a $\uparrow$ symbol.
Methods ML-1M Kuaivideo Book
H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC
Pier 62.74 75.99 65.98 68.74 90.43 45.35 65.11 52.55 53.35 90.93 71.14 80.22 71.62 73.74 92.26
[…]
DNR-G 64.89 77.67 68.0 70.71 92.75 49.61 69.60 57.67 58.39 93.20 74.83 82.57 74.24 77.65 94.46
DNR-B 64.23 77.12 67.30 70.01 92.08 50.30 70.15 58.12 59.16 93.38 75.50 83.53 75.54 78.30 94.23
RQ3: In Table 1, we also observe that the Beta noise alternative generally performs better in Kuaivideo and Book with significant margin, and the Gaussian noise performs better in ML-1M. This indicates the superiority of Beta over Gaussian is not consistent across recommendation platforms. Nevertheless, we observe that the the learned model-based sampler is more aligned with the Beta distribution in practice (see Appendix D.2).
RQ4: To further investigate different model alternatives, we evaluate the different ways of implementing noise injection and the ablation of learning objectives, and present the results in in Table 3. By comparing “w/ G” with “DNR-G or comparing “w/ B” with “DNR-B”, we can see that the full solution with a well optimized model-based sampler consistently outperforms the heuristic sampler.
RQ5: To compare different learning objectives in Eq.(5), we first observe that “w/ G” and “w/B” methods outperforms the Pier baseline, indicating the effectiveness of $\mathcal{L}_{\text{z}}$, which uses the posterior to augment the learning of reranker. Then, all “w/o $\mathcal{L}_{\text{adv}}$” alternatives and “w/o $\mathcal{L}_{\text{x}}$” alternatives appear to downgrade the effectiveness of the solution, compared to the full implementation of DNR. All these observations validate the effectiveness of the derived terms in Eq.(5).
Table 3: Ablation Study for DNR alternatives. “w/ G” and “w/ B” denote heuristic-only posteriors that adopt noise sampler with Gaussian noise and Beta noise without using model-based sampler; “G w/o $\mathcal{L}_{\text{adv}}$” and “B w/o $\mathcal{L}_{\text{adv}}$” represent DNR-G or DNR-B paradigms without adversarial noise learning objective; “G w/o $\mathcal{L}_{\text{x}}$” and “B w/o $\mathcal{L}_{\text{x}}$” refer to DNR-G or DNR-B paradigms without the noise regularization.
Methods ML-1M Kuaivideo Book
H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC
Pier 62.74 75.99 65.98 68.74 90.43 45.35 65.11 52.55 53.35 90.93 71.14 80.22 71.62 73.74 92.26
w/ G 63.36 76.69 66.82 69.34 91.98 47.84 67.34 55.27 56.28 92.43 73.38 81.32 72.89 76.18 93.38
w/ B 63.65 76.90 67.01 69.64 90.85 48.83 68.71 56.60 57.45 92.73 73.42 81.20 72.74 76.23 93.46
G w/o $\mathcal{L}_{\text{adv}}$ 63.71 76.57 66.61 69.43 92.14 48.75 68.70 56.57 57.34 92.21 73.83 81.89 73.27 76.60 93.62
G w/o $\mathcal{L}_{\text{x}}$ 63.47 76.43 66.43 69.19 91.90 48.88 69.10 57.18 57.57 92.69 73.99 82.01 73.44 76.76 94.29
DNR-G 64.89 77.67 68.0 70.71 92.75 49.61 69.60 57.67 58.39 93.20 74.83 82.57 74.24 77.65 94.46
B w/o $\mathcal{L}_{\text{adv}}$ 63.65 76.90 67.01 69.64 90.85 49.55 69.84 58.04 58.31 92.91 74.71 82.49 74.14 77.52 94.01
B w/o $\mathcal{L}_{\text{x}}$ 63.32 76.62 66.76 69.29 90.60 49.96 69.93 58.00 58.77 92.81 74.42 82.20 73.74 77.23 93.96
DNR-B 64.23 77.12 67.30 70.01 92.08 50.30 70.15 58.12 59.16 93.38 75.50 83.53 75.54 78.30 94.23
4.4 Sensitivity Analysis (RQ6)
We conduct sensitivity analysis for $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$ for both DNR-G and DNR-B, and present the results in Figure 3. While the Beta noise generally performs better than Gaussian noise in most settings, there exists some common patterns in these hyper-parameters: 1) all parameters appears to have an optimal value within the searched region, indicating the effectiveness of including each corresponding designs, i.e., $\lambda_{c}$ for the posterior design Eq.(4), $\lambda_{m}$ for the combined reconstruction loss Eq.(6), and $\lambda_{e}$ for the switching from heuristic sampler to model-based sampler; 2) over amplify these hyper-parameters results in the downgrade of the rerank performance, indicating the significance of keeping a balance between the involved components. Additionally, the analysis of $\lambda_{c}$ in the left plot of Figure 3 shows that using the model-based noise sampler can further enhance the performance over the heuristic noise sampler.
Figure 3: Sensitivity analysis of our method on Kuaivideo dataset.
5 Conclusion
In this paper, we provide evidence that the initial retriever scores are informative for the latter reranker stage in the multi-stage recommender system. To fully utilize such scores, we formulate the reranker as a noise reduction problem in the item scoring space and propose DNR, which collaboratively learns a more robust denoising reranker and a generative noise injection posterior. Theoretical analysis and empirical validation across three public datasets demonstrate the effectiveness of the noise reduction formulation and the design of our DNR. We also show that the posterior is intuitively and empirically a beta distribution, which is potentially more suitable than Gaussian noise in recommendation tasks. While a heuristic design of the noisy posterior can effectively improve the model performance, we prove that we can more comprehensively optimize a learnable posterior with the theoretically derived learning objective, and further boost the reranker’s performance.
Appendix A Proofs
A.1 Derivation of the Three Objectives in our Solution
In this section, we illustrate the derivation of the Eq.(5). Similar to the variational Bayesian inference [32], we start by analyzing the KL divergence between the ground truth and estimated posterior:
$D_{\text{KL}}(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u}))$ $=-\mathbb{E}_{p_{\phi}}\Big[\log\frac{p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})}{p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}\Big]$ (9)
$=-\mathbb{E}_{p_{\phi}}\Big[\log\frac{p(\mathbf{x},\mathbf{z}_{u})}{p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}-\log p(\mathbf{z}_{u})\Big]$
$=-\mathbb{E}_{p_{\phi}}\Big[\frac{\log p(\mathbf{x},\mathbf{z}_{u})}{\log p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}\Big]+\log p(\mathbf{z}_{u}),$
Note that a key difference between Eq.(9) and the conventional derivation in [32] is that the variational posterior $\phi$ is not a latent factor encoder that reduces noise and dimensions, but an explicit noise injection model that mimics the behavior of retriever scores. In terms of the denoising problem formulation, it is also similar to a one-step diffusion, but with a key difference in that the ground truth $\mathbf{x}_{u}$ is not sampled by a manually defined distribution. Instead, an existing retriever $p_{\text{x}}$ serves as the prior.
Then, rearranging the terms of Eq.(9), we can derive the corresponding objectives that influence the data likelihood as follows:
$-\log p(\mathbf{z}_{u})$ (10)
$=$ $-\mathbb{E}_{p_{\phi}}\big[\log\frac{p(\mathbf{x},\mathbf{z}_{u})}{p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}\big]-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
$=$ $-\mathbb{E}_{p_{\phi}}\big[\log p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})\big]+D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\big)-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
$=$ $-\mathbb{E}_{p_{\phi}}\big[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})\big]+\mathbb{E}_{p_{\phi}}\big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})}\big]+D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\big)$
$-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
which is equivalent to Eq.(5). It generally follows the derivation of the evidence lower bound, but we further decompose the reconstruction term into the parameterized reconstruction loss $\mathcal{L}_{\text{z}}$ and the adversarial posterior learning loss $\mathcal{L}_{\text{adv}}$. This decomposition indicates that the reranker $q_{\theta}$ is simultaneously an approximation of the ground truth $p_{\text{z|x}}$ as well as a premise factor that determines $p_{\text{z|x}}$, and their difference mainly describes the reranker’s error in predicting user feedback.
A.2 The Probabilistic View of Standard Reranking Models
In this section, we formally analyze the Eq.(3). Recall that the direct optimization goal $\mathcal{L}_{\text{std}}=-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})]$, where the observed $\mathbf{x}_{u}$ only comes from the retriever prior $p_{\text{x}}$ without posterior modeling. Then, we can extract the error of $q_{\theta}$ and derive the following relation:
$-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u})]$ $=-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}\Big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x}_{u})}\Big]-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}[\log p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x}_{u})]$ (11)
$=-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}\Big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x}_{u})}\Big]+D_{\text{KL}}\big(p_{\text{x}}(\mathbf{x}_{u})\|p_{\text{x|z}}(\mathbf{x}_{u}|\mathbf{z}_{u})\big)-\log p(\mathbf{z}_{u})$
$\Rightarrow-\log p(\mathbf{z}_{u})=$ $-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u})]+\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}\Big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x}_{u})}\Big]$
$-D_{\text{KL}}\big(p_{x}(\mathbf{x}_{u})\|p_{\text{x|z}}(\mathbf{x}_{u}|\mathbf{z}_{u})\big).$
which corresponds to Eq.(3).
As we have illustrated in section 3.2, the direct optimization of the first term $\mathcal{L}_{\text{direct}}$ might be misaligned with the data likelihood maximization goal when there exists irregular retriever scores or the posterior is inconsistent with the prior distribution. In contrast, our proposed solution solves this limitation by collaboratively learning the noise-injecting posterior $p_{\phi}$ along with the reranker. The resulting objectives in Eq.(5) intentionally finds a posterior that align with the ground truth prior and explore adversarial cases that the reranker fail to accurately reconstruct the user feedback (details in section 3.4).
Appendix B Learning Algorithm
We present the overall learning process of our method in Algorithm 1.
0: Hyperparameters $\lambda_{c}$ (or $\lambda_{e}$ if using learnable posterior), $\lambda_{m}$
0: Optimized denoising model $q_{\theta}$ (and the posterior $p_{\phi}$ if using learnable posterior).
1:for Epoch $i$ do
2:  if Using learnable posterior and $i>\lambda_{e}$ then
3:   for Each sample $(\mathbf{u},\mathbf{x}_{u},\mathbf{z}_{u})$ do
4:    Sample $\mathbf{x}\sim p_{\phi}^{\text{model}}$ for given $\mathbf{u}$ and $\mathbf{z}_{u}$;
5:    Optimize $\theta$ based on $\mathcal{L}_{\theta}$;
6:    Optimize $\phi$ based on $\mathcal{L}_{\text{adv}}$ and $\mathcal{L}_{\text{x}}^{\prime}$;
7:   end for
8:  else
9:   for Each sample $(\mathbf{u},\mathbf{x}_{u},\mathbf{z}_{u})$ do
10:    Sample $\mathbf{x}\sim p_{\phi}^{\text{heuristic}}$ for given $\mathbf{u}$ and $\mathbf{z}_{u}$
11:   end for
12:  end if
13:end for
Algorithm 1 Training procedures of DNR
Appendix C Detailed Settings for Offline Experiments
C.1 Datasets
The ML-1M dataset is a public benchmark in the field of recommender systems, comprising 1 million ratings from 6,040 users across more than 3,900 movies. The Kuaivideo dataset is sourced from Kuaishou, a popular short-video app with over 300 million active users daily. The Amazon Books dataset is a comprehensive collection of product reviews specifically focused on books available on Amazon. For all datasets, we initially process by eliminating users and items with fewer than 20 interactions, to avoid the cold-start issue. For the retriever stage, we split the data into train and test subsets at a ratio of 8:2. For the reranker stage, we sort the interactions in chronological order, using the last six interactions as the item list revealed to users after reranking. We present the statistics of the processed datasets for the retriever and reranker stages in Table LABEL:dataset_stage1 and Table LABEL:dataset_stage2, respectively.
Table 4: Dataset Statistics for Retriever Stage
Dataset # Users # Items # Actions
ML-1M 6,020 3,043 995,154
Kuaivideo 89,310 10,395 3,270,132
Amazon-Toys 35,732 38,121 1,960,674
Table 5: Dataset Statistics for Reranker Stage
Dataset # Users # Items # Sequences
ML-1M 6,022 3,043 161,646
Kuaivideo 89,416 10,395 513,010
Amazon-Toys 35,736 38,121 311,386
C.2 Baselines
We detail the compared baselines of our main experiments in the following, including traditional recommenders, list-refinement methods, generator-evaluator methods, and diffusion-based methods. Traditional Recommenders: predict the scores of candidate items and rank them accordingly.
  • SASRec [23] proposes a self-attention based sequential recommendation model (SASRec) that captures long-term semantics and short-term dynamics using an attention mechanism.
  • Caser [36] introduces a convolutional sequence embedding recommendation model that captures sequential patterns using convolutional filters applied to latent space embeddings.
  • GRU4Rec [37] apply recurrent neural networks (RNNs) to session-based recommendations, addressing the challenge of short user sessions without long-term profiles.
  • MiDNN [24] proposes a global optimization framework for e-commerce search ranking that considers mutual influences between items, using extended features and sequence generation to optimize ranking and maximize overall utility.
List-Refinement Methods: encode the item lists from the previous stage and refine the rankings based on this information.
  • DLCM [8] leverages local ranking context from top-retrieved documents to refine initial ranking lists. The model uses a recurrent neural network to capture document interactions and an attention-based loss function, significantly improving performance over traditional learning-to-rank methods.
  • SetRank [38] introduces a neural learning-to-rank model that learns a permutation-invariant ranking function directly on document sets. Utilizing multi-head self-attention blocks, SetRank captures cross-document interactions and achieves robust performance across varying input sizes.
  • PRM [6] addresses personalized re-ranking in recommendation systems by incorporating user-specific preferences into the re-ranking process. It considers user behavior and candidate list, enhancing both relevance and personalization in the final recommendations.
  • MIR [25] proposes to capture complex interactions between user actions and candidate list features across multiple levels, improving the accuracy and relevance of recommendations through a hierarchical reranking approach.
Generator-Evaluator Methods: generate multiple ranked item lists and evaluate them to select the best one for users.
  • EGRerank [22] proposes an evaluator-generator framework for learning-to-rank (LTR) in e-commerce applications. The framework includes an evaluator that evaluates recommendations involving context, a generator that maximizes the evaluator score through reinforcement learning, and a discriminator that ensures the generalization of the evaluator.
  • Pier [9] follows a two-stage architecture with a Fine-grained Permutation Selection Module (FPSM) and an Omnidirectional Context-aware Prediction Module (OCPM). FPSM selects top-K candidate permutations based on user interest using SimHash, while OCPM evaluates these permutations with an omnidirectional attention mechanism.
  • NAR4Rec [10] explores the use of non-autoregressive generative models for re-ranking in recommendation systems, addressing challenges such as sparse training samples and dynamic candidates by introducing a matching model, unlikelihood training to distinguish feasible sequences, and contrastive decoding to capture item dependencies.
Diffusion-Based Methods: utilize diffusion models to generate items or lists for recommendation.
  • DiffuRec [26] incorporates a diffusion process to inject noise into target item embeddings and a reverse process to reconstruct the target item representation.
  • DCDR [11] introduces a new framework that leverages diffusion models for the reranking stage in recommendation systems. DCDR extends traditional diffusion models with a discrete forward process and a conditional reverse process to generate high-quality item sequences.
C.3 Implementation of Diverse Strategies for Leveraging Retriever Scores
Here, we present the detailed implementation of different ways to leverage retriever scores in Figure C.3. The left one refers to “c score”, which leverages a concatenate operation to combine the score features with encoded user states from users’ interaction history. The moderate one represents “+ score”, which incorporates the score feature with user states by a simple plus operation. The right one is the “w score”, which adopts the initial candidate score as the weight of candidate items.s
Figure 4: Different ways to leverage retriever scores.
Appendix D Complementary Experimental Analysis
D.1 Experiments of Retriever Stage
Detailed Experimental Settings for Retriever Stage
We implement the retriever stage with the MF method [39], which aims to improve the click rate performance. The dimension of the item embedding is 128, and the learning rate is 0.001. We randomly select negative items to train the MF method, enhancing the retriever’s performance despite the item sparsity.
Experimental results of Retriever Stage
Here, we present the detailed results of the retriever stage in Table 6.
Table 6: Performance of the retriever stage, which adopts the matrix factorization (MF) model to rate and select the top-50 items from full item set.
Datasets ML-1M Kuaivideo Book
AUC 79.20% 89.62% 88.82%
D.2 Visualization of Learned Noises
To visualize the injected noise distribution, we compare the density curve of the injected noise from different noise injection strategies, including Gaussian noise, Beta noise, and learning-based noise. As demonstrated in Figure 5, the distribution of learned noise is more aligned with the distribution of true noise between retriever scores and user feedback. This means that the resulting posterior with model-based noise sampler better reduces the divergence term $\delta_{\text{x}}$ in Eq.(5), and potentially improve the alignment between our proposed solution with the three objectives and the data log-likelihood maximization goal.
Figure 5: The visualization of injected noise from different variants.
D.3 Detailed Results of Sensitivity Analysis
Here, we present the detailed results of the sensitivity analysis for different datasets in Figure 6, Figure 7, and Figure 8.
Figure 6: Sensitivity of DNR to the hyperparameter $\lambda_{c}$ on different datasets, which controls the degree of noise injection.
Figure 7: Sensitivity of DNR to the hyperparameter $\lambda_{e}$ on different datasets, which decides the epochs where the adversarial learning starts.
Figure 8: Sensitivity of DNR to the hyperparameter $\lambda_{m}$ on different datasets, which controls the magnitude of the augmentation.
D.4 Online A/B Test
To better investigate the effectiveness of our DNR in an online environment, we conduct an A/B test on an industrial platform with video feeds to 100+ millions of users daily.
Detailed Experimental Settings for Online Test The multi-stage pipeline of online system consists of three stages: retrieval, ranking, and reranking. In the final reranking stage, the top-6 items for user exposure are selected from a candidate set of 50 items generated by the ranking stage. During the ranking process, a scoring model computes multiple dimensional features for each item. Among these, we utilize PCTR(predicted probability that the video will be watched for more than 3 seconds) as the primary retrieval score $\mathbf{x}_{u}$. DNR is applied to generate a new set of scores for 50 selected items and the top-6 items are chosen as the final exposure .
Experiment Results For our online experiments, the total traffic is randomly divided, with 12.5% allocated to the DNR method and 12.5% to the baseline method. Each experiment is conducted for at least 7 days to ensure the reliability of the results. The DNR method demonstrates a significant improvement in realshow(realshow refers to the cumulative number of videos watched by users.), increasing it by 1.089% . These results indicate that the DNR method can estimate PCTR more accurately, which is strongly correlated with realshow.
Appendix E Limitation and Discussion
E.1 Limitations
Computational Cost: Compared to a direct optimization framework with Eq.(2) that only learns $\theta$, our framework requires extra efforts to train a posterior model $\phi$ (if using learnable posterior) with three additional loss terms to back-propagate. This might multiply the training cost for several folds. Fortunately, $\phi$ is not involved in inference, so the inference cost stays the same.
Finding Correct Distribution Models: In this paper, we present an implementation based on the Beta distribution assumption of the prior and the posterior. As we have mentioned in section 3.3, finding the optimal setting for the heuristic posterior might require expert experience, and we are skeptical on whether there exists a fixed setting for all user requests.
E.2 Future Directions
Multi-task Scenarios: When user can engage with an item with multiple behaviors (e.g., like and share a news story), DNR may adapt to this scenario by simply formulating a multi-dimensional response space for $\mathbf{z}_{u}$ and $\mathbf{x}_{u}$. Yet, different behavior signals may typically follow a heterogeneous distribution where a uniform modeling technique might become sub-optimal. Thus, we consider it beneficial to further investigate how the proposed framework may accommodate the multi-task solutions.
Better Approximation for KL Divergence: In our solution, the noise regularization term $\mathcal{L}_{\text{x}}$ requires the estimation of the distribution $p_{\phi}$ and $p_{\text{x}}$. We approximate this by directly aligning the output scores with $\mathcal{L}_{\text{x}}$ in Eq.(8) in the implementation to avoid estimation of the prior. To this end, there might exist better solutions to approach the original $\mathcal{L}_{\text{x}}$ term.
More Investigation on Beta Posterior: We have illustrated that the ground truth posterior and the prior of retriever scores are likely to be a Beta distribution. Yet, there still lack analytical results on the optimal personalized parameters of $\alpha$ and $\beta$ that are aligned with the ground truth. Further study on these parameters may reveal user behavioral patterns in the real world, which could further help improve the recommendation performance.

### v2

@@
-Robust Denoising Neural Reranker for Recommender Systems
+Denoising Neural Reranker for Recommender Systems
 Abstract
-For multi-stage recommenders in industry, a user request would first trigger a simple and efficient retriever module that selects and ranks a list of relevant items, then calls a slower but more sophisticated deep reranking model that refines the item arrangement before exposure to the user.
-The latter model typically reranks the item list conditioned on the users’ history content and the initial ranking from retrievers.
-Although this two-stage retrieval-ranking framework demonstrates practical effectiveness, the significance of retriever scores from the previous stage has been limitedly explored, which is informative.
-In this work, we first theoretically analyze the limitations of using retriever scores as the reranker’s input directly and argue that the reranking task is essentially a noise reduction problem from the retriever scores.
-Following this notion, we derive an adversarial framework, DNR, that associates the denoising reranker with a carefully designed noise generation module.
-We extend the conventional score error minimization term with three augmented objectives, including: 1) a denoising objective that aims to denoise the noisy retriever scores to align with the user feedback; 2) an adversarial retriever score generation objective that improves the exploration in the retriever score space; and 3) a distribution regularization term that aims to align the distribution of generated noisy retriever scores with the real ones.
-Extensive experiments are conducted on three public datasets, together with analytical support, validating the effectiveness of the proposed DNR.
+For multi-stage recommenders in industry, a user request would first trigger a simple and efficient retriever module that selects and ranks a list of relevant items, then the recommender calls a slower but more sophisticated reranking model that refines the item list exposure to the user.
+To consistently optimize the two-stage retrieval reranking framework, most efforts have focused on learning reranker-aware retrievers.
+In contrast, there has been limited work on how to achieve a retriever-aware reranker.
+In this work, we provide evidence that the retriever scores from the previous stage are informative signals that have been underexplored.
+Specifically, we first empirically show that the reranking task under the two-stage framework is naturally a noise reduction problem on the retriever scores, and theoretically show the limitations of naive utilization techniques of the retriever scores.
+Following this notion, we derive an adversarial framework DNR that associates the denoising reranker with a carefully designed noise generation module.
+The resulting DNR solution extends the conventional score error minimization loss with three augmented objectives, including: 1) a denoising objective that aims to denoise the noisy retriever scores to align with the user feedback; 2) an adversarial retriever score generation objective that improves the exploration in the retriever score space; and 3) a distribution regularization term that aims to align the distribution of generated noisy retriever scores with the real ones.
+We conduct extensive experiments on three public datasets and an industrial recommender system, together with analytical support, to validate the effectiveness of the proposed DNR.
 1 Introduction
-Recommender systems aim to personalize a list of items that maximizes user engagement signals such as purchases, video plays, and link sharing.
-To achieve this, industrial solutions [1, 2, 3] typically employ a multi-stage architecture comprising a retriever and a reranker.
-The retriever, which may further decompose into several consecutive sub-stages of retrieval and ranking models, efficiently filters a large item pool down to a manageable candidate set of relevant items.
+Recommender systems aim to personalize a list of items that maximizes user engagement signals, such as purchases in e-commerce, video plays in video platforms, and link sharing in social networks.
+To achieve a feasible solution with large-scale candidate pools and user requests, industrial solutions Covington et al.
+(2016); Ricci et al.
+(2022); Liu et al.
+(2025) typically employ a multi-stage architecture comprising a retriever and a reranker.
+The retriever, which may further decompose into several more sub-stages of retrieval and ranking modules, efficiently filters a large item pool down to a manageable candidate set of relevant items.
 Then the reranker adopts a more sophisticated but more accurate model that refines the candidate item set into an optimal ordering for the list-wise exposure.
-Formally, the goal of the latter reranking task [4, 5, 6, 7] is to learn an ordering function over the retrieved candidates that maximizes user responses on the top-reranked items.
-This task is particularly challenging due to the combinatorial nature of the item permutation space.
-Exhaustively searching for the optimal list exposure in this space becomes infeasible as the candidate set grows, making this a key bottleneck in real-world recommender systems.
-To overcome this challenge, early studies propose to refine the initial retriever scores [5, 8, 6] by modeling the item mutual influences in the candidate item list.
-Recent state-of-the-art solutions find it more effective to frame the reranking task as a generative modeling task [9, 10, 11] where the list generation probability should directly align with the user’s list-wise rewards.
-Yet, these methods largely overlook the potential of initial retriever scores, which may offer rich prior information to enhance the reranking performance.
+To optimize the two-stage framework towards a consistent goal of recommendation (e.g., aligning true user behavior or preference) across all stages, the majority of efforts have been focusing on learning a reranker-aware retriever Gallagher et al.
+(2019); Xu et al.
+(2024a).
+However, there is limited work on how to achieve a retriever-aware reranker.
+Formally, the goal of the latter reranking task Carbonell & Goldstein (1998); Jin et al.
+(2008); Pei et al.
+(2019); Liu et al.
+(2022) is to learn an ordering function over the retrieved candidates that maximizes user responses on the top-reranked items.
+Early studies propose to re-score the initial retrieval items Jin et al.
+(2008); Ai et al.
+(2018); Pei et al.
+(2019) by modeling the mutual influences among items in the candidate list.
+Recent state-of-the-art approaches find it more effective to formulate the reranking task as a list generation problem Shi et al.
+(2023); Ren et al.
+(2024); Lin et al.
+(2024), where the lists with better list-wise rewards are given higher generation probability.
+Yet, existing methods largely overlook the potential of initial retriever scores in the input space, which may offer rich prior information to enhance the reranking performance.
 In practice, a straightforward approach to leverage initial retriever scores is to include them as input features directly, which shows promise to improve reranking performance as in Figure 1(a).
-Reranking as a noise reduction problem: Though effective, this naive solution neglects the noisy nature of the initial retriever scores, which may cause model robustness concerns and downgrade the effectiveness of the learned reranking model.
-Specifically, the retriever stage, constrained by strict latency and computational budgets, usually adopts a simpler model than the reranker.
-In contrast, the reranker, operating on a much smaller candidate set, can use a more sophisticated architecture and thus achieves significantly higher accuracy (Figure 1(b)).
-In other words, by assuming higher noise in the retriever scores, it naturally indicates a noise reduction process of the reranking stage.
-This noise reduction pattern is more evident when we observe the changes in the embedding representations of ground truth target items, comparing Figure 1(c) and Figure 1(d).
-This means that the reranker behaves like a denoising model even without explicit modeling of the retriever’s noise.
-However, as we will show in section 3.1, the retriever’s noise theoretically contributes to the misalignment between the traditional objective of reranking and the user response likelihood maximization goal, restricting the generalization ability of the reranker.
-As a countermeasure, we need a more robust solution that can learn a denoising reranker that is aware of the retriever’s noisy score distribution, while taking advantage of the information from retriever scores.
+Still, as we will illustrate in section 3.2, this naive solution might be far from exploiting this retriever information effectively to align with user feedback, which is addressed by the following contributions in this work.
+Formulating reranking as a noise reduction problem: The retriever stage typically employs a simpler model than the reranker to process large candidate pools, given computational budget constraints.
+In contrast, the reranker, operating on a much smaller candidate set, can use a more sophisticated architecture and thus achieves significantly higher accuracy (Figure 1(a), 1(b)).
+What follows is an empirical assumption that there is higher noise in the retriever scores than the reranker scores, which naturally indicates a noise reduction process of the reranking stage.
+This noise reduction pattern is more evident when we observe the changes in the embedding representations of ground truth target items.
+We show evidence by comparing Figure 1(c) and Figure 1(d), where the same set of retrieved candidates is not distinguished well by the retriever, but the target items are better predicted by the reranker, in terms of the embedding space.
+This further proves that the reranker behaves like a denoising model on the retriever’s noise.
+However, directly aligning the user behavior on the condition of the observed retriever score, ignoring the uncertainty of this noisy signal, may cause a misalignment with the user behavior alignment goal in terms of the entire system.
+Thus, we argue that learning a noise-aware reranker is crucial for achieving a retriever-aware reranker.
 (a) Initial retriever scores further boost the performance.
@@
 The green shaded circle in (d) represents the reranker’s selection of the items of exposure.
-The proposed solution: Based on the previous analysis, we propose to augment the denoising neural reranker (DNR) with a carefully designed noise injection model, which estimates the initial retriever score’s posterior distribution.
-We theoretically show that the data likelihood maximization goal derives a solution framework that includes three learning objectives: 1) an augmented denoise objective that aims to boost the user feedback reconstruction under unseen retriever scores, increasing the reranker’s effectiveness to denoise noisy retriever scores; 2) an adversarial retriever score generation objective that aims to generate irregular retriever scores for the reranker that are hard to denoise, which in turn improves the reranker’s robustness; and 3) a score posterior regularization term that aims to align the noise-aware posterior model with the initial retriever score distribution, guaranteeing data alignment with the system behavior in previous stages.
-Additionally, we show that the noise-injecting posterior and the denoising reranker collaborate with each other in a similar way to the one-step diffusion framework [12], but standard Gaussian noise is not necessarily an empirically optimal choice.
-Instead, we illustrate that the binomial nature of the user response better suits a Beta distribution in the retriever score space, which derives our heuristic solution for the posterior.
-Still, this naive design lacks personalization abilities.
-To mitigate this limitation, we design a learnable posterior model that approximately optimizes the three augmented learning objectives.
-Empirical supports and findings: We conduct empirical experiments on three public datasets to verify the superiority of DNR against leading rerankers.
-Notably, the reranker paired with the learnable posterior best aligns with the observed ground-truth posterior, which theoretically improves the alignment between the reranker’s learning objective and the user feedback data likelihood maximization goal, resulting in the best empirical reranking performance in our experiment.
-In general, this work provides evidence that the initial retriever scores may be more informative than we think, and a more comprehensive utilization of this information may derive a solution that surpasses state-of-the-art methods.
+The proposed solution: Motivated by the previous analysis, we propose an adversarial framework, which formulates a denoising neural reranker (DNR) with a carefully designed noise generation module, which introduces unseen noise to users’ feedback to augment the distribution modeling of retriever scores.
+We theoretically show that the user behavior alignment goal can be decomposed into three learning objectives in our solution framework: 1) a denoising objective that aims to denoise the noisy retriever scores under both observed and synthetic retriever scores; 2) an adversarial objective that encourages the noise generator to synthesize adversarial samples that are hard to denoise, which in turn improves the reranker’s effectiveness; and 3) a score distribution regularization term that aims to align the distribution between synthetic retriever scores with the real retriever scores.
+We conduct empirical experiments on three public datasets and an industrial recommender system to verify the superiority of DNR against naive utilization strategies of retriever scores, as well as leading reranker models.
+Additionally, we find that the three proposed learning objectives in the adversarial denoising framework yield better recommendation performance than a single denoising objective.
 2 Related Work
 2.1 Multi-stage Recommendation
-Modern content-intensive web services with large candidate pools typically employ multi-stage recommender systems [13, 14, 7] (also known as cascade ranking [15, 16] process) to balance accuracy and efficiency under strict latency constraints.
+Modern content-intensive web services with large candidate pools typically employ multi-stage recommender systems Qin et al.
+(2022); Zheng et al.
+(2024); Liu et al.
+(2022) (also known as cascade ranking Liu et al.
+(2017); Gallagher et al.
+(2019) process) to balance accuracy and efficiency under strict latency constraints.
 Each stage progressively narrows down the candidate set while employing increasingly sophisticated models.
-While most existing work focuses on independently optimizing individual stages (e.g., retrieval [17, 18] or reranking stage [7, 6]), recent studies explore joint optimization across the entire pipeline [16, 13, 19].
-These approaches leverage later-stage models to guide earlier-stage learning, improving overall coherence and performance [20, 21].
+While most existing work focuses on independently optimizing individual stages (e.g., retrieval Wang et al.
+(2011; 2017) or reranking stage Liu et al.
+(2022); Pei et al.
+(2019)), recent studies explore joint optimization across the entire pipeline Gallagher et al.
+(2019); Qin et al.
+(2022); Xu et al.
+(2024b).
+These approaches leverage later-stage models to guide earlier-stage learning, improving overall coherence and performance Zhao et al.
+(2023); Xu et al.
+(2024a).
 However, our work takes a complementary direction—–using earlier-stage scores to regulate later-stage optimization.
 2.2 Reranking Strategies
-In sequential recommendation, re-ranking [6, 22, 8] aims to refine the initial ranking list generated by a previous retrieval stage.
+In sequential recommendation, re-ranking Pei et al.
+(2019); Huzhang et al.
+(2023); Ai et al.
+(2018) aims to refine the initial ranking list generated by a previous retrieval stage.
 Existing re-ranking approaches can be broadly categorized into the following paradigms.
-Single-point re-scoring methods [23, 24] independently predict a refined score for each item and re-rank them accordingly.
+Single-point re-scoring methods Kang & McAuley (2018); Zhuang et al.
+(2018) independently predict a refined score for each item and re-rank them accordingly.
 While efficient, they ignore dependencies between items, potentially leading to suboptimal list-level performance.
-Unlike single-point approaches, list-refinement methods [6, 8, 25] explicitly model mutual influences between items by treating the pre-ranked candidate list as input.
-Techniques such as Transformer-based architectures [6] capture user preference to optimize the list holistically.
-Generator-evaluator methods [22, 9] first generate multiple candidate lists and then evaluate them using a learned utility function to select the best one.
+Unlike single-point approaches, list-refinement methods Pei et al.
+(2019); Ai et al.
+(2018); Xi et al.
+(2022) explicitly model mutual influences between items by treating the pre-ranked candidate list as input.
+Techniques such as Transformer-based architectures Pei et al.
+(2019) capture user preference to optimize the list holistically.
+Generator-evaluator methods Huzhang et al.
+(2023); Shi et al.
+(2023) first generate multiple candidate lists and then evaluate them using a learned utility function to select the best one.
 They can effective generate a high-quality recommendation list but suffer from the increasing computational cost due to the generation-evaluation loop.
-Emerging techniques leverage diffusion models [11, 26] to generate refined item lists by iteratively denoising a ranking distribution, which excel in capturing complex user preferences.
+Emerging techniques leverage diffusion models Lin et al.
+(2024); Li et al.
+(2024) to generate refined item lists by iteratively denoising a ranking distribution, which excel in capturing complex user preferences.
 2.3 Adversarial Learning
-Adversarial learning [27] has emerged as a powerful technique for data augmentation in recommendation systems, improving the models robustness and generalization.
-By leveraging generative adversarial networks (GANs) [18] or adversarial training [28, 29], these methods synthesize high-quality user-item interactions or perturb existing data to enhance model robustness.
-For instance, IRGAN [18] employs a minimax game between a generator (sampling hard negative items) and a discriminator (distinguishing real vs.
+Adversarial learning Lowd & Meek (2005) has emerged as a powerful technique for data augmentation in recommendation systems, improving the models robustness and generalization.
+By leveraging generative adversarial networks (GANs) Wang et al.
+(2017) or adversarial training He et al.
+(2018); Tang et al.
+(2020), these methods synthesize high-quality user-item interactions or perturb existing data to enhance model robustness.
+For instance, IRGAN Wang et al.
+(2017) employs a minimax game between a generator (sampling hard negative items) and a discriminator (distinguishing real vs.
 synthetic interactions) to improve ranking performance.
-Similarly, AdvIR [30] introduces adversarial perturbations to embedding spaces, forcing the model to learn more generalizable representations.
+Similarly, AdvIR Park & Chang (2019) introduces adversarial perturbations to embedding spaces, forcing the model to learn more generalizable representations.
 These methods demonstrate that adversarial learning can not only augment training data but also enhance models’ robustness.
-Our method is designed to add noise by adversarial learning, which can enhance the denoise models (rerankers)’ robustness to uncertain noise in the pre-ranking scores.
+Our method is designed to add noise by adversarial learning, which can enhance the denoising models (rerankers)’ effectiveness in denosing the pre-ranking scores to align with user feedback.
 3 Methodology
@@
 3.1 Problem Formulation
-For a given user request $\mathbf{u}$ (which may contain information about the candidate items $\mathcal{I}(u)$, user profile features, user interaction history), we observe the following during data collection: The candidate set $\mathcal{I}(u)=\{i_{1},i_{2},\cdots,i_{n}\}$ retrieved by the retriever; The corresponding (continuous) retriever scores of $\mathcal{I}(u)$ are $\mathbf{x}_{u}=[x_{i_{1}},x_{i_{2}},\cdots,x_{i_{n}}]\in[0,1]^{n}$, which essentially describes the use feedback likelihood in the view of the retriever; The ground truth user feedback (i.e., binary labels such as click, watch, share, etc.) of the candidate set, i.e., $\mathbf{z}_{u}=[z_{i_{1}},z_{i_{2}},\cdots,z_{i_{n}}]\in\{0,1\}^{n}$.
-Note that for recommender systems that further select K items after the reranking process, we may not observe the user feedback for the remaining items.
-Yet, from a probabilistic viewpoint, we can assume zeros as defaults since the item passing the reranking process is the premise of observing the user feedback signals.
-The Noise Reduction Task: In our formulation, we denote the initial retriever as the prior $p_{\text{x}}(\mathbf{x}|\mathbf{u})$ and the learnable reranker $q_{\theta}(\mathbf{z}|\mathbf{x},\mathbf{u})$ as the conditional user feedback likelihood estimator .
-The $q_{\theta}$ predicts the user feedback given the retriever scores and the user context, i.e., $\hat{\mathbf{z}}_{u}\sim q_{\theta}(\mathbf{z}|\mathbf{x}_{u},\mathbf{u})$, and the learning goal adopts the general data likelihood maximization objective:
-$\max p(\mathbf{z}_{u}|\mathbf{u})$ (1)
-where the ground-truth $p(\mathbf{z}_{u}|\mathbf{u})$ is determined by the retriever $p_{\text{x}}$, the reranker $q_{\theta}$, and the user.
-Note: For simplicity of the presentation, we omit the user condition term $\mathbf{u}$ for the remainder of the paper (e.g.
-we will denote the reranker as $q_{\theta}(\mathbf{z}|\mathbf{x}_{u})$ and simplify the learning goal as $\max p(\mathbf{z}_{u})$), since we adopt an analysis framework on the user request level.
+For a given user request $\mathbf{u}$ (which may contain information about the candidate items $\mathcal{I}(u)=\{i_{1},i_{2},\cdots,i_{n}\}$ retrieved by the retriever, user profile features, and user interaction history), we observe the following during data collection: 1) The corresponding (continuous) retriever scores of $\mathcal{I}(u)$ are $\mathbf{x}_{u}=[x_{i_{1}},x_{i_{2}},\cdots,x_{i_{n}}]\in[0,1]^{n}$, which essentially describe the user preference prediction in the view of the retriever; 2) The ground truth user feedback (i.e., binary labels such as click, watch, share, etc.) of the candidate set, i.e., $\mathbf{z}_{u}=[z_{i_{1}},z_{i_{2}},\cdots,z_{i_{n}}]\in\{0,1\}^{n}$.
+Note that for recommender systems that further select K items after the reranking process, we may never observe the positive user feedback for the remaining nKn-K items.
+From a probabilistic viewpoint, we can assume zero labels as defaults for these items, since passing the reranking process is the premise of observing the user feedback signals.
+The Noise Reduction Task: In our formulation, the overall recommendation-feedback process $\mathbf{u}\rightarrow\mathbf{x}\rightarrow\mathbf{z}$ (where $\mathbf{x}$ and $\mathbf{z}$ are random variables for retriever scores and user feedback) involves an initial retriever $p_{\text{x}}(\mathbf{x}|\mathbf{u})$, which serves as the retriever score generation prior (and is assumed non-optimizable in the reranking task).
+The learnable reranker $q_{\theta}(\mathbf{z}|\mathbf{x},\mathbf{u})$ serves as the conditional user feedback likelihood estimator that predicts the user feedback $\hat{\mathbf{z}}$ given the retriever scores $\mathbf{x}$ and the user context $\mathbf{u}$, i.e., $\hat{\mathbf{z}}\sim q_{\theta}(\cdot|\mathbf{x}_{u},\mathbf{u})$.
+Given an observed feedback $\mathbf{z}_{u}$, the learning goal adopts the general data log likelihood maximization objective:
+$\max_{\theta}\log p(\mathbf{z}_{u}|\mathbf{u},\theta)$ (1)
+where the predicted likelihood $p(\mathbf{z}_{u}|\mathbf{u},\theta)$ for the observed user feedback $\mathbf{z}_{u}$ is determined by the retriever $p_{\text{x}}$, the reranker $q_{\theta}$, and the user request $\mathbf{u}$.
+Note: For simplicity of the presentation, we omit the user condition term $\mathbf{u}$ for the remainder of the paper (e.g., we will denote the retriever as $p_{\text{x}}(\mathbf{x})$, the reranker as $q_{\theta}(\mathbf{z}|\mathbf{x}_{u})$, and simplify the learning goal as $\max_{\theta}\log p(\mathbf{z}_{u})$), since we adopt an analysis framework on the user request level.
 3.2 Analytical Limitation of Direct Optimization Methods
@@
 which maximizes the user behavior alignment conditioned on the given retrieval scores.
-In practice, according to the binary nature of $\mathbf{z}_{u}$, we can simplify the learning objective as standard binary cross-entropy (BCE) loss.
-Theoretically, we can find that $\mathcal{L}_{\text{direct}}$ optimizes an upper bound of the negative data log likelihood (details in Appendix A.2):
+As a special case, according to the binary nature of $\mathbf{z}_{u}$, we can simplify the learning objective as standard binary cross-entropy (BCE) loss.
+Theoretically, we can find that $\mathcal{L}_{\text{direct}}$ optimizes an upper bound of the negative data log likelihood (details in Appendix B.2):
 $-\log p(\mathbf{z}_{u})$ $=\mathcal{L}_{\text{direct}}+L_{1}+L_{2}$ (3)
@@
 $L_{2}$ $=-D_{\text{KL}}\big(p_{\text{x}}(\mathbf{x})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big).$
-where $p_{\text{z|x}}$ represents the ground truth user feedback probability conditioned on retriever scores, and $p_{\text{x|z}}$ represents the ground truth posterior.
+where $p_{\text{z|x}}$ represents the ground-truth user feedback probability and $p_{\text{x|z}}$ represents the posterior distribution of retriever scores conditioned on user feedback, both terms are not related to the reranker $q_{\theta}$ and they are merely determined by the retriever prior and the user.
 Intuitively, smaller values of $L_{1}$ and $L_{2}$ make the solution of Eq.(2) more aligned with the goal of Eq.(1).
-However, these two terms are not fully controlled by the direct optimization framework and may generate a non-robust solution of $q_{\theta}$:
-Irregular cases from retriever: The expectation term $L_{1}$ describes the difference between $q_{\theta}$ and the ground truth $p_{\text{z|x}}$ under the retriever prior.
-This term is non-positive and is minimized when $q_{\theta}(\mathbf{z}|\mathbf{x})=p(\mathbf{z}|\mathbf{x})$ for any variables.
-Though $L_{1}$ is expected to be small when the reranker’s predictions align with the user feedback in dominant cases, the retriever may happen to generate an $\mathbf{x}_{u}$ that the reranker fails to comprehend, inducing a significant divergence between the optimization of Eq.(2) and the learning goal of Eq.(1).
-Regular cases with noisy retriever scores: The KL term $L_{2}$ describes the difference between the retriever score distribution $p_{\text{x}}$ and the obtained posterior given the knowledge of $\mathbf{z}_{u}$.
-This term is also non-positive and can be non-trivial due to the noisy nature of the retriever (as illustrated in Figure 1).
-This means that there might exist consistently significant misalignment between Eq.(2) and Eq.(1).
-3.3 Implementing Noise Injection Posterior
-To overcome the aforementioned limitations, we propose to complement the denoising reranker $q_{\theta}$ with a generative posterior $p_{\phi}(\mathbf{x}|\mathbf{z}_{u})$ that explicitly models the ground truth retriever score posterior distribution $p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})$.
-As described in Section 1, $p_{\text{x|z}}$ behaves as a noise-injecting model, so:
-$\mathbf{x}_{u}^{\prime}=(1-\lambda_{c})\mathbf{z}_{u}+\lambda_{c}\bm{\epsilon}_{\phi},$ (4)
-where the sampled error term $\bm{\epsilon}_{\phi}$ represents the noisy behavior of the posterior. To this end, we investigate two choices of noise samplers for $\bm{\epsilon}_{\phi}$:
-  • Heuristic Sampler: This option uses a non-parametric sampler $\bm{\epsilon}_{\phi}^{\text{heuristic}}$ with a manually designed distribution.
-For example, we can adopt standard Gaussian noises similar to [31, 26], but we notice that this is not a good representation of the ground truth.
-Intuitively, given the binary user response signals, a natural assumption of the retriever prior is the Beta distribution, where the parameters depend on the user request $\mathbf{u}$.
-What follows is a Beta posterior distribution, which derives our heuristic solution as $\bm{\epsilon}_{\phi}\sim\text{Beta}(\alpha,\beta)$.
-For simplicity, we set $\alpha=\beta=0.5$ to approximate the uncertainty, which would help improve the chance of generating harder samples for the reranker and consequently improve the robustness of the solution.
-Note that this implementation assumes that the noise does not depend on the user, and the personalization only comes from the label $\mathbf{z}_{u}$;
-  • Model-based Sampler: While the heuristic sampler is effective in our empirical study, we are skeptical on whether the fixed setting of $\alpha$ and $\beta$ is optimal for all user requests.
-To achieve a personalized noise estimation, we use a learnable model $\bm{\epsilon}_{\phi}^{\text{model}}$ with parameter $\phi$ that will be optimized along with the reranker (details in the next section).
-As we will show in Section 4.3, this choice is empirically better, and the learned error exhibits a Beta distribution without explicitly defining it.
-Without specification, our standard DNR solution uses this model-based sampler.
+However, these two terms are not fully controlled by the direct optimization framework, potentially causing significant misalignment between Eq.(2) and Eq.(1).
+This will lead to a discrepancy between the final reranking results and user feedback, and thus, the suboptimal recommendation performance.
+3.3 Implementing Noise Generator for the Posterior
+To overcome the aforementioned limitations, we propose to complement the denoising reranker $q_{\theta}$ with a noise generator $f_{\phi}(\cdot|\mathbf{z}_{u})$.
+Formally, we generate noise with $\bm{\epsilon}\sim f_{\phi}$ and add it to the user feedback to form the synthetic retriever score posterior $p_{\phi}(\mathbf{x}|\mathbf{z}_{u})$, taking insight from the reparameterization trick Kingma & Welling (2014):
+$\mathbf{x}_{u}^{\prime}=(1-\lambda_{c})\mathbf{z}_{u}+\lambda_{c}\bm{\epsilon},$ (4)
+where the generated noise $\bm{\epsilon}$ represents the noisy behavior in retriever scores $\mathbf{x}_{u}^{\prime}$, $\lambda_{c}$ represents the proportion of noise in the generated retriever scores.
+We denote the generation process as $\mathbf{x}_{u}^{\prime}\sim p_{\phi}$, which is analogous to the sampling process with ground truth posterior (assumed unknown and intractable), i.e., $\mathbf{x}\sim p_{\text{x|z}}$.
+For implementation, we investigate two choices of noise generator $f_{\phi}$:
+  • Heuristic Generator: This option uses a non-parametric noise generator $f_{\phi}^{\text{heuristic}}$ with a manually designed distribution.
+For example, we can adopt random Gaussian noise similar to Ho et al.
+(2020); Li et al.
+(2024), but we notice that this is not a good representation of the noise in retriever scores.
+Intuitively, given the binary user feedback signals, a natural assumption of the retriever prior is the Beta distribution, where we can derive the heuristic solution as $\bm{\epsilon}\sim\text{Beta}(\alpha,\beta)$.
+For simplicity, one may set $\alpha=\beta=0.5$ to approximate the uncertainty.
+In practice, we find that it is also beneficial to set $\alpha$ and $\beta$ that align with the observed noise distribution in the data.
+  • Model-based Generator: While the heuristic generator is effective in our empirical study, we are skeptical of whether the heuristic noise generator is optimal for augmenting the retriever score distribution.
+To achieve a more precise estimation, we design a learnable model $f_{\phi}^{\text{model}}$, with parameter $\phi$, that will be optimized along with the reranker (details in the next section).
+The generated noise $\bm{\epsilon}$ should also be aligned with the observed distribution, and it also participates in an adversarial training objective in order to better explore the retriever score space, which in turn improves the reranker’s effectiveness.
+As we will show in Section 4.3, this choice is empirically better than the heuristic generator.
 3.4 Overall Learning Framework
-In this section, we show how to optimize the user feedback alignment i.e., Eq.(1) by collaborating the learning of the reranker and the discovery of better noise-injecting posterior.
-Specifically, following a variational inference framework, we can decompose the negative log-likelihood objective into three loss terms (detailed derivation in Appendix A.1):
-$-\log p(\mathbf{z}_{u})$ $=\mathcal{L}_{\text{z}}+\mathcal{L}_{\text{x}}+\mathcal{L}_{\text{adv}}+\delta_{\text{x}}$ (5)
+In this section, we show how to optimize the user feedback alignment goal i.e., Eq.(1) by simultaneously learning the denoising reranker and the adversarial noise generator.
+Specifically, we can decompose the negative log-likelihood objective into three loss terms (derivation in Appendix B.1):
+$-\log p(\mathbf{z}_{u})$ $=\mathcal{L}_{\text{z}}+\mathcal{L}_{\text{adv}}+\mathcal{L}_{\text{x}}+\delta_{\text{x}}$ (5)
 $\mathcal{L}_{\text{z}}$ $=-\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\Big[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})\Big]$
@@
 $\mathcal{L}_{\text{x}}$ $=D_{\text{KL}}\Big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\Big),$
-which consists of an augmented feedback reconstruction loss term $\mathcal{L}_{\text{z}}$, an adversarial noise exploration loss $\mathcal{L}_{\text{adv}}$, and a posterior distribution regularization term $\mathcal{L}_{\text{x}}$.
-The additional term $\delta_{\text{x}}=-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)\leq 0$ represents the divergence between the noise injection posterior $p_{\phi}$ and the corresponding ground truth $p_{\text{x|z}}$.
-Intuitively, we expect to achieve a small divergence in $\delta_{\text{x}}$ and $L_{\text{x}}$ so that the optimization of the three terms is aligned with the goal of Eq.(1).
-Denoising with Augmented Retriever Scores: $\mathcal{L}_{\text{z}}$ in Eq.(5) denotes the reconstruction loss under the noisy retriever scores sampled by the posterior.
-Specifically, we can first sample $\mathbf{x}_{u}^{\prime}\sim p_{\phi}(\cdot|\mathbf{z}_{u})$ as the generated retriever scores, then pass them into the reranker $q_{\theta}$ for the reconstruction of $\mathbf{z}_{u}$.
-The optimization can also be simplified into the BCE loss with $\mathbf{z}_{u}$ as labels, and it only updates $\theta$.
+which consists of an augmented denoising loss term $\mathcal{L}_{\text{z}}$, an adversarial noise generation loss $\mathcal{L}_{\text{adv}}$, and a score distribution regularization term $\mathcal{L}_{\text{x}}$.
+We will illustrate the details of these three terms in the following sections.
+The additional term $\delta_{\text{x}}=-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)\leq 0$ represents the negative divergence between the generated noisy retriever scores $p_{\phi}$ and the corresponding real $p_{\text{x|z}}$.
+This term is not directly optimizable due to the unknown $p_{\text{x|z}}$, but we expect to achieve a small divergence in $\delta_{\text{x}}$ with the optimization of the other three terms, so that the overall optimization is aligned with the goal of Eq.(1).
+Denoising with Augmented Retriever Scores: $\mathcal{L}_{\text{z}}$ in Eq.(5) denotes the denoising loss under the noisy retriever scores generated by the noise generator.
+Specifically, we can first synthesize $\mathbf{x}_{u}^{\prime}\sim p_{\phi}(\cdot|\mathbf{z}_{u})$ as the generated retriever scores according to Eq.(4), then pass them into the reranker $q_{\theta}$ for the prediction of user feedback $\mathbf{z}_{u}$, and $\mathcal{L}_{\text{z}}$ only updates $\theta$.
+The optimization can also be simplified into the BCE loss with $\mathbf{z}_{u}$ as labels.
 Note that the generated scores $\mathbf{x}_{u}^{\prime}$, different from the observed $\mathbf{x}_{u}$, essentially serve as the augmented scores that the initial retriever would have produced given the user request.
-This means that the reranker will have to learn a more robust reconstruction model of the ground-truth user feedback.
-Combining with the direct optimization, we have the final reconstruction loss for the reranker:
+This means that the reranker will be aware of the noisy behavior of the retriever.
+Combining with the direct optimization, we have the final denoising loss for the reranker:
 $\mathcal{L}_{\theta}=\mathcal{L}_{\text{direct}}+\lambda_{m}\mathcal{L}_{\text{z}},$ (6)
-where $\lambda_{m}$ controls the magnitude of the augmentation, and the posterior $p_{\phi}$ is fixed.
-Adversarial Noise Learning: $\mathcal{L}_{\text{adv}}$ describes how the reranker behaves differently from the true user feedback under the viewpoint of the noise-injecting posterior $p_{\phi}$.
+where $\lambda_{m}$ controls the magnitude of the augmentation, and the parameters of the noise generator $f_{\phi}$ are fixed.
+Additionally, our model-agnostic framework does not restrict the architecture of the denoising reranker $q_{\theta}$, as we will show in section 4.3, one can adopt various backbones, including PRM Pei et al.
+(2019) and PIER Shi et al.
+(2023).
+Adversarial Noise Learning: $\mathcal{L}_{\text{adv}}$ describes how the reranker $q_{\theta}$ behaves differently from the user feedback under the viewpoint of the synthetic retriever scores.
 In practice, this term is expected to be small since the ratio $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})\approx 1$ in major cases due to the user feedback alignment of Eq.(6) and the augmented objective $\mathcal{L}_{\text{z}}$.
-Nevertheless, in the view of the posterior, we can see that $\mathcal{L}_{\text{adv}}$ encourages adversarial samples of $\mathbf{x}$ that produce small ratios of $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$.
-In other words, a better $p_{\phi}$ should be able to find irregular samples that are correctly recovered by the true distribution $p_{\text{z|x}}$ but are incorrectly reconstructed by the reranker $q_{\theta}$.
-For the posterior with heuristic sampler $\bm{\epsilon}_{\phi}^{\text{heuristic}}$, it is not optimizable once $\alpha$ and $\beta$ are chosen.
-In contrast, for the model-based sampler $\bm{\epsilon}_{\phi}^{\text{model}}$, we follow the intuition of irregular case exploration, and approximate $\mathcal{L}_{\text{adv}}$ by first fixing the reranker and then optimizing the posterior with a reconstruction minimization loss:
+Nevertheless, in the view of generated noisy retriever scores, we can see that $\mathcal{L}_{\text{adv}}$ encourages adversarial samples of $\mathbf{x}$ that produce small ratios of $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$.
+In other words, a better $p_{\phi}$ should be able to explore and synthesize retriever scores that are incorrectly denoised by the reranker $q_{\theta}$.
+For the heuristic noise generator $f_{\phi}^{\text{heuristic}}$, it is not optimizable once $\alpha$ and $\beta$ are chosen.
+In contrast, for the model-based generator $f_{\phi}^{\text{model}}$, we follow the intuition of adversarial case exploration, and approximate $\mathcal{L}_{\text{adv}}$ by first fixing the reranker and then optimizing the generator with an adversarial loss:
 $\min_{\phi}\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u}^{\prime}).$ (7)
-In our design, we directly output $\mathbf{x}_{u}^{\prime}$ using $\phi$ so that the end-to-end learning becomes feasible for Eq.(7).
-In cases where the sampling process is not directly optimizable, one may have to consider other alternatives like reparameterization tricks [32] or policy gradient [33].
-On the other hand, $\mathcal{L}_{\text{adv}}$ does not encourage over-exploration of $\mathbf{x}_{u}^{\prime}$, which may negatively impact the learning of the reranker.
-In such cases, the generated scores are unrealistically different from the true posterior, diverging from the user behavior, inducing small ground truth $p_{\text{z|x}}$ and large ratio of $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$, violating $\mathcal{L}_{\text{adv}}$.
-Noise Regularization: $\mathcal{L}_{\text{x}}$ estimates the distributional divergence between the posterior estimator $p_{\phi}$ and the prior $p_{\text{x}}$, and we aim to find a $p_{\phi}$ that reduces this divergence.
-Intuitively, this term favors the posteriors that can mimic the noisy behavior of the prior $p_{\text{x}}$ and remove the additional accuracy-related information from $\mathbf{z}$.
-We can calculate this term by first sampling retriever scores as $\mathbf{x}^{\prime}_{u}\sim p_{\phi}(\cdot|\mathbf{z}_{u})$, then estimating their log ratio difference from that of the prior.
-In cases where we do not have the model to directly estimate $p_{\text{x}}(\mathbf{x}_{u}^{\prime})$, one can circumvent this issue by directly aligning $\mathbf{x}_{u}^{\prime}$ with the observed samples $\mathbf{x}_{u}$:
-$\mathcal{L}_{\text{x}}^{\prime}\approx-\mathbb{E}_{p_{\phi}}\big[\log\frac{\mathbf{x}_{u}}{\mathbf{x}^{\prime}_{u}}\big],$ (8)
-which is an instance-level objective more rigorous than $\mathcal{L}_{\text{x}}$.
-Intuitively, a smaller value of $\mathcal{L}_{\text{x}}^{\prime}$ means that $\phi$’s output is close to the prior’s output $\mathbf{x}_{u}$, which restricts the distribution-level objective $\mathcal{L}_{\text{x}}$ to small values.
-Yet, we remind readers that it is not the case vice versa.
-Additionally, it is infeasible to repeat the exact same request and obtain augmented samples to cover the prior distribution of $p_{\text{x}}$.
-Fortunately, the collaborative learning across different user requests will approximate the sampling process of pair $(\mathbf{x}^{\prime}_{u},\mathbf{x}_{u})$ with similar patterns.
-Again, note that the heuristic sampler $\bm{\epsilon}_{\phi}^{\text{heuristic}}$ is not optimizable by this objective, but one can achieve a small value of $\mathcal{L}_{\text{x}}$ by carefully designing the distribution (e.g., choosing Beta instead of Gaussian).
-For the training of the posterior with model-based sampler $\bm{\epsilon}_{\phi}^{\text{model}}$, we first adopt the heuristic sampler in the first $\lambda_{e}$ epochs to ensure learning stability, then engage the adversarial learning of $\mathcal{L}_{\text{adv}}$ and the optimization of Eq.(8) afterwards, switching the sampler to $\bm{\epsilon}_{\phi}^{\text{model}}$.
-As the summary of our solution, we present our training procedure in Appendix B.
+We implement the model-based generator $f_{\phi}^{\text{model}}$ by a 2-layer MLP and directly output $\mathbf{x}_{u}^{\prime}$ using Eq.(4), so that the end-to-end learning becomes feasible for Eq.(7).
+In cases where the noise generator is not directly optimizable, one may have to consider other alternatives like reparameterization tricks Kingma & Welling (2014) or policy gradient Silver et al.
+(2014).
+Notably, $\mathcal{L}_{\text{adv}}$ does not encourage over-exploration of $\mathbf{x}_{u}^{\prime}$, which may negatively impact the learning of the reranker.
+In such cases, the synthetic scores are unrealistically different from the real retriever scores, diverging from the user feedback, inducing small $p_{\text{z|x}}$ and large $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$, violating $\mathcal{L}_{\text{adv}}$.
+Noise Regularization: $\mathcal{L}_{\text{x}}$ aims to reduce the divergence between the synthetic retriever score distribution $p_{\phi}$ and the real retriever score prior distribution $p_{\text{x}}$.
+Intuitively, this term favors the synthetic score distribution $p_{\phi}$ that can mimic the noisy behavior of the real retriever’s scores $p_{\text{x}}$.
+For implementation, we first synthesize the retriever score $\mathbf{x}_{u}^{\prime}$ with Eq.(4) and calculate the distributional difference between synthetic scores from $p_{\phi}$ and retriever scores from $p_{\text{x}}$.
+Then, we minimize this difference to approximate the distribution-level KL divergence.
+Again, given that the heuristic noise generator $f_{\phi}^{\text{heuristic}}$ is not optimizable by the objectives in Eq (5), it might not be the optimal choice for augmenting the retriever score distribution.
+As a consequence, this could lead to recommendations that are not in alignment with user feedback.
+For the training of the model-based noise generator $f_{\phi}^{\text{model}}$, we first adopt the heuristic generator in the first $\lambda_{e}$ epochs to ensure learning stability of reranker $q_{\theta}$ throughout $\mathcal{L}_{z}$, then engage the adversarial learning of $\mathcal{L}_{\text{adv}}$ and the optimization of $\mathcal{L}_{\text{x}}$ afterward, switching the generation of $\bm{\epsilon}$ from heuristic generator to model-based generator.
+We present the algorithm of our training procedure in Appendix C.
 4 Experiments
 In this section, we conduct extensive experiments to answer the following questions:
-RQ1: How does DNR perform in the reranking tasks of recommendation systems compared with leading baselines?
+RQ1: How does DNR perform in the reranking tasks of recommendation systems compared with leading rerankers?
 RQ2: How well is the denoising formulation compared to alternative methods of leveraging retriever scores?
-RQ3: What is the performance difference between Gaussian noise and Beta noise?
-RQ4: How does the model-based sampler compare to the heuristic sampler in practice?
+RQ3: Can the denoising formulation generalize to other reranker backbones, rather than PRM?
+RQ4: What is the performance difference between the heuristic noise generator and the adversarial model-based generator?
 RQ5: How does different objectives in DNR contribute to the performance?
@@
 Datasets.
-We conduct experiments on three real-world datasets, including ML-1M [34], Kuaivideo [3], and Amazon-books [35].
-The detailed description and the statistics of the processed datasets for both the retriever and the reranker stages are presented in Appendix C.1.
-Baselines: We compare the performance of DNR with four categories of recommendation methods in our reranking task, including traditional recommenders (i.e., SASRec [23], Caser [36], GRU4Rec [37], and MiDNN [24]), which individually predict the scores of candidate items and rank them accordingly; list-refinement methods (i.e., DLCM [8], SetRank [38], PRM [6], and MIR [25]), which encode the mutual inference in candidate items and refine the rankings based on this information; generator-evaluator methods (i.e., EGRerank [22], Pier [9], and NAR4Rec [10]), which generate multiple reranked item lists and evaluate them to select the best one for users; and diffusion-Based methods (i.e., DiffuRec [26] and DCDR [11]) which utilize diffusion models to generate items or lists for recommendation.
-Detailed information for our baselines is presented in Appendix C.2.
+We conduct experiments on three real-world datasets, including ML-1M Harper & Konstan (2016), Kuaivideo Liu et al.
+(2025), and Amazon-books Ni et al.
+(2019).
+The detailed description and the statistics of the processed datasets for both the retriever and the reranker stages are presented in Appendix D.1.
+For the online experimental verification, please refer to Appendix E.6.
+Baselines: We compare the performance of DNR with four categories of recommendation methods in our reranking task, including traditional recommenders (i.e., SASRec Kang & McAuley (2018), Caser Tang & Wang (2018), GRU4Rec Hidasi et al.
+(2016), and MiDNN Zhuang et al.
+(2018)), which individually predict the scores of candidate items and rank them accordingly; list-refinement methods (i.e., DLCM Ai et al.
+(2018), SetRank Pang et al.
+(2020), PRM Pei et al.
+(2019), and MIR Xi et al.
+(2022)), which encode the mutual inference in candidate items and refine the rankings based on this information; generator-evaluator methods (i.e., EGRerank Huzhang et al.
+(2023), Pier Shi et al.
+(2023), and NAR4Rec Ren et al.
+(2024)), which generate multiple reranked item lists and evaluate them to select the best one for users; and diffusion-Based methods (i.e., DiffuRec Li et al.
+(2024) and DCDR Lin et al.
+(2024)) which utilize diffusion models to generate items or lists for recommendation.
+Detailed information for our baselines is presented in Appendix D.3.
 Implementation Details: Our experiments are implemented on Nvidia A40 GPUs with Python 3.8 and PyTorch 1.12.
-We construct the retriever stage with the collaborative filtering method [39] to obtain the initial retriever scores for the top-50 candidate items.
-The detailed setting and experimental results of the retriever stage are presented in Appendix D.1.
+We construct the retriever stage with the collaborative filtering method Koren et al.
+(2009) to obtain the initial retriever scores for the top-50 candidate items.
+The detailed setting and experimental results of the retriever stage are presented in Appendix E.1.
 The reranker reorders these items filtered by retrievers and selects the top $K=6$ for display.
@@
 We tune the hyperparameters with $\lambda_{c}\in[0.1,1.0]$, $\lambda_{m}\in[0.1,1.0]$, and $\lambda_{e}\in[0,200]$.
+The optimal hyperpapemters adopted for our experimental results are presented in Appendix D.2.
 Evaluation Metrics: Following common practices, we employ Hit Ratio@6 (H@6), NDCG@6 (N@6), MAP@6 (M@6), F1@6, and AUC for performance evaluation of the reranker stage.
@@
 We compare the performance of our proposed DNR with multiple leading baselines and the results are detailed in Table 1.
-“DNR-G” and “DNR-B” denote two alternatives to implement $\bm{\epsilon}_{\phi}^{\text{heuristic}}$ by injecting Gaussian noise and Beta noise, respectively, before switching to the model-based sampler.
-We can observe that the generator-evaluator-based methods (i.e., EGRerank [22], Pier [9], and NAR4Rec [10]) achieve comparable performance among the multiple baselines, validating the effectiveness of the generator-evaluator paradigm in the reranker stage.
+“DNR-G” and “DNR-B” denote using two alternatives of $f_{\phi}^{\text{heuristic}}$ to generate Gaussian noise and Beta noise, respectively, before switching to the model-based noise generator $f_{\phi}^{\text{model}}$ in the epoch $\lambda_{e}$.
+We can observe that the generator-evaluator-based methods (i.e., EGRerank Huzhang et al.
+(2023), Pier Shi et al.
+(2023), and NAR4Rec Ren et al.
+(2024)) achieve comparable performance among the multiple baselines, validating the effectiveness of the generator-evaluator paradigm in the reranker stage.
 In comparison, our DNR solutions consistently outperform the state-of-the-art baselines across the three benchmarks.
-This highlights the effectiveness of our noise reduction formulation and the significance of improving reranker robustness in our solution.
+This highlights the superiority of our noise reduction formulation and the significance of improving reranker effectiveness with three objectives in adversarial learning.
 Table 1: Overall performance of different methods for the reranking recommendation.
@@
 4.3 Ablation Study (RQ2, RQ3, RQ4, & RQ5)
-RQ2: To investigate different ways to leverage retriever scores, we develop three PRM variants for the noise generator: “c score” refers to leveraging a concatenate operation to combine the score features with encoded user states from users’ interaction history; “+ score” denotes incorporating the score feature with user states by a simple addition; “w score” represents adopting the initial retriever scores as the weights of candidate items.
-The frameworks of the three variants are shown in Appendix C.3.
+RQ2 & RQ3: To investigate different ways to leverage retriever scores, we develop three variants for leveraging retriever scores as rerankers’ input: “c score” refers to leveraging a concatenate operation to combine the score features with encoded user states from users’ interaction history; “+ score” denotes incorporating the score feature with user states by a simple addition; “w score” represents adopting the initial retriever scores as the weights of candidate items.
+The frameworks of the three variants are shown in Appendix D.4.
+To validate the generalizability of DNR across different reranker backbones, we conduct experiments comparing our three variants and DNR itself when using PRM Pei et al.
+(2019) and Pier Shi et al.
+(2023) as the underlying rerankers, respectively.
 We report the experimental results in Table 2.
-The variants “c score”, “+ score”, and “w score” achieve better performance than the PRM backbone and slightly better performance than the state-of-the-art baseline Pier, validating that it is beneficial to include the information from retriever scores in the reranking stage.
-Additionally, our DNR solutions outperform all three variants, indicating a superior strategy to leverage the retriever scores.
-Table 2: Ablation Study for the different ways of leveraging the retriever scores. Performance that exceeds Pier’s is indicated with a $\uparrow$ symbol.
+The variants “c score”, “+ score”, and “w score” all achieve superior performance compared to the base reranker backbone, confirming that incorporating retriever scores into the reranking stage yields tangible benefits.
+Moreover, our DNR outperforms all three variants, underscoring denoising formulation as a more effective strategy for leveraging retriever scores.
+Importantly, DNR delivers consistent improvements across both backbones, thereby validating the generalizability of our denoising formulation for reranking architectures.
+Table 2: Ablation Study for the different ways of leveraging the retriever scores on both PRM and Pier reranker backbones.
 Methods ML-1M Kuaivideo Book
 H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC
+PRM 60.09 72.85 62.21 65.51 88.20 39.92 55.93 42.97 46.18 85.15 67.86 76.88 66.44 70.42 92.00
+[…]
+[…]
+[…]
+DNR 64.23 77.12 67.30 70.01 92.08 50.30 70.15 58.12 59.16 93.38 75.50 83.53 75.54 78.30 94.23
 Pier 62.74 75.99 65.98 68.74 90.43 45.35 65.11 52.55 53.35 90.93 71.14 80.22 71.62 73.74 92.26
@@
 […]
-DNR-G 64.89 77.67 68.0 70.71 92.75 49.61 69.60 57.67 58.39 93.20 74.83 82.57 74.24 77.65 94.46
-DNR-B 64.23 77.12 67.30 70.01 92.08 50.30 70.15 58.12 59.16 93.38 75.50 83.53 75.54 78.30 94.23
-RQ3: In Table 1, we also observe that the Beta noise alternative generally performs better in Kuaivideo and Book with significant margin, and the Gaussian noise performs better in ML-1M.
-This indicates the superiority of Beta over Gaussian is not consistent across recommendation platforms.
-Nevertheless, we observe that the the learned model-based sampler is more aligned with the Beta distribution in practice (see Appendix D.2).
-RQ4: To further investigate different model alternatives, we evaluate the different ways of implementing noise injection and the ablation of learning objectives, and present the results in in Table 3.
-By comparing “w/ G” with “DNR-G or comparing “w/ B” with “DNR-B”, we can see that the full solution with a well optimized model-based sampler consistently outperforms the heuristic sampler.
-RQ5: To compare different learning objectives in Eq.(5), we first observe that “w/ G” and “w/B” methods outperforms the Pier baseline, indicating the effectiveness of $\mathcal{L}_{\text{z}}$, which uses the posterior to augment the learning of reranker.
-Then, all “w/o $\mathcal{L}_{\text{adv}}$” alternatives and “w/o $\mathcal{L}_{\text{x}}$” alternatives appear to downgrade the effectiveness of the solution, compared to the full implementation of DNR.
+DNR 63.41 77.28 68.02 70.22 91.97 48.89 69.35 57.76 56.02 92.38 73.68 82.61 74.82 76.88 93.78
+RQ4: To further investigate different noise generators, we conduct ablation experiments on the heuristic noise generators that generate noise from the Gaussian distribution and Beta distribution, denoted as “w/ G” and “w/ B” respectively.
+The results are presented in Table 3.
+By comparing “w/ G” with “DNR-G or comparing “w/ B” with “DNR-B”, we can see that our DNR framework—equipped with a model-based noise generator for adversarial learning—consistently outperforms counterparts using heuristic noise generators.
+RQ5: To compare different learning objectives in Eq.(5), we first observe that “w/ G” and “w/B” methods outperform the PRM baseline, indicating the effectiveness of $\mathcal{L}_{\text{z}}$, which uses the heuristic noise generator to augment the learning of reranker.
+Then, all “w/o $\mathcal{L}_{\text{adv}}$” alternatives and “w/o $\mathcal{L}_{\text{x}}$” alternatives appear to downgrade the effectiveness of the full implementation of DNR.
 All these observations validate the effectiveness of the derived terms in Eq.(5).
 Table 3: Ablation Study for DNR alternatives.
-“w/ G” and “w/ B” denote heuristic-only posteriors that adopt noise sampler with Gaussian noise and Beta noise without using model-based sampler; “G w/o $\mathcal{L}_{\text{adv}}$” and “B w/o $\mathcal{L}_{\text{adv}}$” represent DNR-G or DNR-B paradigms without adversarial noise learning objective; “G w/o $\mathcal{L}_{\text{x}}$” and “B w/o $\mathcal{L}_{\text{x}}$” refer to DNR-G or DNR-B paradigms without the noise regularization.
+“G w/o $\mathcal{L}_{\text{adv}}$” and “B w/o $\mathcal{L}_{\text{adv}}$” represent DNR-G or DNR-B paradigms without adversarial noise learning objective; “G w/o $\mathcal{L}_{\text{x}}$” and “B w/o $\mathcal{L}_{\text{x}}$” refer to DNR-G or DNR-B paradigms without the noise regularization.
 Methods ML-1M Kuaivideo Book
 H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC H@6 N@6 M@6 F1@6 AUC
-Pier 62.74 75.99 65.98 68.74 90.43 45.35 65.11 52.55 53.35 90.93 71.14 80.22 71.62 73.74 92.26
+PRM 60.09 72.85 62.21 65.51 88.20 39.92 55.93 42.97 46.18 85.15 67.86 76.88 66.44 70.42 92.00
 w/ G 63.36 76.69 66.82 69.34 91.98 47.84 67.34 55.27 56.28 92.43 73.38 81.32 72.89 76.18 93.38
@@
 We conduct sensitivity analysis for $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$ for both DNR-G and DNR-B, and present the results in Figure 3.
-While the Beta noise generally performs better than Gaussian noise in most settings, there exists some common patterns in these hyper-parameters: 1) all parameters appears to have an optimal value within the searched region, indicating the effectiveness of including each corresponding designs, i.e., $\lambda_{c}$ for the posterior design Eq.(4), $\lambda_{m}$ for the combined reconstruction loss Eq.(6), and $\lambda_{e}$ for the switching from heuristic sampler to model-based sampler; 2) over amplify these hyper-parameters results in the downgrade of the rerank performance, indicating the significance of keeping a balance between the involved components.
-Additionally, the analysis of $\lambda_{c}$ in the left plot of Figure 3 shows that using the model-based noise sampler can further enhance the performance over the heuristic noise sampler.
-Figure 3: Sensitivity analysis of our method on Kuaivideo dataset.
+While the Beta noise generally performs better than Gaussian noise in most settings, as shown in Table 2, there exist some common patterns in these hyper-parameters: 1) all parameters appears to have an optimal value within the searched region, indicating the effectiveness of including each corresponding designs, i.e., $\lambda_{c}$ for the synthetic noisy scores in Eq.(4), $\lambda_{m}$ for the combined denoising loss Eq.(6), and $\lambda_{e}$ for the switching from heuristic noise generator to model-based one; 2) Over-amplifying these hyper-parameters results in the downgrade of the rerank performance, indicating the significance of keeping a balance between the involved components.
+Additionally, the analysis of $\lambda_{c}$ in the left plot of Figure 3 shows that using the model-based noise generator can further enhance the performance over the heuristic one.
+Figure 3: Sensitivity analysis of hyperparameters (i.e., $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$) for our method on the Kuaivideo dataset.
 5 Conclusion
 In this paper, we provide evidence that the initial retriever scores are informative for the latter reranker stage in the multi-stage recommender system.
-To fully utilize such scores, we formulate the reranker as a noise reduction problem in the item scoring space and propose DNR, which collaboratively learns a more robust denoising reranker and a generative noise injection posterior.
+To formulate the relationship between retriever and reranker scores, we view the reranker as a noise reduction task from the retriever scores and propose DNR.
+To fully explore the score space, we propose the framework of DNR, adding noise to generate retriever scores as augmentation and then denoising them to align with user feedback.
+While a heuristic design of the noise distribution can effectively improve the model performance (e.g., Gaussian or Beta distribution), we prove that adversarially learns a denoising reranker and a noise generator for the retriever stage can further boost the reranker’s performance.
 Theoretical analysis and empirical validation across three public datasets demonstrate the effectiveness of the noise reduction formulation and the design of our DNR.
-We also show that the posterior is intuitively and empirically a beta distribution, which is potentially more suitable than Gaussian noise in recommendation tasks.
-While a heuristic design of the noisy posterior can effectively improve the model performance, we prove that we can more comprehensively optimize a learnable posterior with the theoretically derived learning objective, and further boost the reranker’s performance.
-Appendix A Proofs
-A.1 Derivation of the Three Objectives in our Solution
+Ethics Statement.
+This work is designed to explore the significance of retriever scores for the reranker, formulating the reranker as a denoising process of retriever scores with an adversarial framework.
+We do not foresee any direct, immediate, or negative societal impacts of our research.
+Reproducibility Statement.
+All the results in this work are reproducible. We have discussed the optimal hyperparameters and the details on devices and software environments in Section 4.1 and Appendix D.2.
+Appendix A Notations
+symbol description
+$\mathbf{u}$ user request information including profile features and interaction history
+$\mathbf{x}$ random variables of retriever scores on the candidate item set
+$\mathbf{z}$ random variables of user feedback on the candidate item set
+$\mathbf{x}_{u}$, $\mathbf{z}_{u}$ observed retriever scores and user feedback in data
+$\mathbf{x}_{u}^{\prime}$ the generated retriever scores from the learned or heuristic posterior model $\phi$
+$q_{\theta}$ the probabilistic representation of the reranker model
+$p_{\phi}$ the probabilistic representation of the posterior
+$f_{\phi}$ the noise generator that implements the posterior
+$p_{\text{x}}$ the probabilistic representation of the non-optimizable retriever
+$p_{\text{x|z}}$ the posterior of a given reranker
+$p_{\text{z|x}}$ the probabilistic representation of ground truth user feedback
+$\mathcal{L}_{\text{direct}}$ the standard user feedback alignment loss given the retriever scores
+$\mathcal{L}_{\text{z}}$ the user feedback alignment loss given generated retriever scores from posterior
+$\mathcal{L}_{\text{adv}}$ the adversarial loss that encourages the posterior to generate irregular noises
+$\mathcal{L}_{\text{x}}$ the regularization of posterior
+$\lambda_{c}$ the hyperparameter that controls the magnitude of noise injection in Eq.(4)
+$\lambda_{m}$ the hyperparameter that balances $\mathcal{L}_{\text{direct}}$ and $\mathcal{L}_{z}$ when learning the reranker
+$\lambda_{e}$ the hyperparameter that determines which epoch switches the noise generator
+Table 4: List of Notations
+Appendix B Proofs
+B.1 Derivation of the Three Objectives in our Solution
 In this section, we illustrate the derivation of the Eq.(5).
-Similar to the variational Bayesian inference [32], we start by analyzing the KL divergence between the ground truth and estimated posterior:
+Similar to the variational Bayesian inference Kingma & Welling (2014), we start by analyzing the KL divergence between the real retriever scores and synthetic noisy scores:
 $D_{\text{KL}}(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u}))$ $=-\mathbb{E}_{p_{\phi}}\Big[\log\frac{p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})}{p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}\Big]$ (8)
@@
 $=-\mathbb{E}_{p_{\phi}}\Big[\frac{\log p(\mathbf{x},\mathbf{z}_{u})}{\log p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}\Big]+\log p(\mathbf{z}_{u}),$
-Note that a key difference between Eq.(9) and the conventional derivation in [32] is that the variational posterior $\phi$ is not a latent factor encoder that reduces noise and dimensions, but an explicit noise injection model that mimics the behavior of retriever scores.
-In terms of the denoising problem formulation, it is also similar to a one-step diffusion, but with a key difference in that the ground truth $\mathbf{x}_{u}$ is not sampled by a manually defined distribution.
-Instead, an existing retriever $p_{\text{x}}$ serves as the prior.
+Note that a key difference between Eq.(8) and the conventional derivation in Kingma & Welling (2014) is that the variational posterior $\phi$ is not a latent factor encoder that reduces noise and dimensions, but an explicit noise generator that mimics the behavior of retriever scores.
+In terms of the denoising problem formulation, it is also similar to a one-step diffusion, but with a key difference in that the real retriever scores $\mathbf{x}_{u}$ are not sampled by a manually defined distribution.
+Instead, an existing retriever $p_{\text{x}}$ serves as the prior distribution.
 Then, rearranging the terms of Eq.(8), we can derive the corresponding objectives that influence the data likelihood as follows:
 $-\log p(\mathbf{z}_{u})$ (9)
-$=$ $-\mathbb{E}_{p_{\phi}}\big[\log\frac{p(\mathbf{x},\mathbf{z}_{u})}{p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}\big]-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
-$=$ $-\mathbb{E}_{p_{\phi}}\big[\log p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})\big]+D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\big)-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
-$=$ $-\mathbb{E}_{p_{\phi}}\big[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})\big]+\mathbb{E}_{p_{\phi}}\big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})}\big]+D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\big)$
+$=$ $-\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log\frac{p(\mathbf{x},\mathbf{z}_{u})}{p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}\big]-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
+$=$ $-\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})\big]+D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\big)-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
+$=$ $-\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})\big]+\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})}\big]+D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x}}(\mathbf{x})\big)$
 $-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)$
 which is equivalent to Eq.(5).
-It generally follows the derivation of the evidence lower bound, but we further decompose the reconstruction term into the parameterized reconstruction loss $\mathcal{L}_{\text{z}}$ and the adversarial posterior learning loss $\mathcal{L}_{\text{adv}}$.
-This decomposition indicates that the reranker $q_{\theta}$ is simultaneously an approximation of the ground truth $p_{\text{z|x}}$ as well as a premise factor that determines $p_{\text{z|x}}$, and their difference mainly describes the reranker’s error in predicting user feedback.
-A.2 The Probabilistic View of Standard Reranking Models
+It generally follows the derivation of the evidence lower bound in variational auto-encoders, except that we further decompose the first reconstruction term into the augmented denoising loss $\mathcal{L}_{\text{z}}=-\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})\big]$ and the adversarial learning loss $\mathcal{L}_{\text{adv}}=\mathbb{E}_{\mathbf{x}\sim p_{\phi}}\big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})}\big]$.
+Intuitively, when $q_{\theta}$ is closely aligned with the real user feedback distribution $p_{\text{z|x}}$, the adversarial term $\mathcal{L}_{\text{adv}}$ will be close to zero.
+Minimizing $\mathcal{L}_{\text{adv}}$ would encourage the noise generator $f_{\phi}$ to find samples where $q_{\theta}$ and $p_{\text{z|x}}$ behave differently.
+B.2 The Analytical Limitation of Standard Reranking Models
 In this section, we formally analyze the Eq.(3).
-Recall that the direct optimization goal $\mathcal{L}_{\text{std}}=-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x})]$, where the observed $\mathbf{x}_{u}$ only comes from the retriever prior $p_{\text{x}}$ without posterior modeling.
-Then, we can extract the error of $q_{\theta}$ and derive the following relation:
+Recall that the direct optimization goal $\mathcal{L}_{\text{direct}}|_{\mathbf{x}=\mathbf{x}_{u}}=-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u})]$, where the observed $\mathbf{x}_{u}$ only comes from the retriever prior $p_{\text{x}}$ without posterior modeling.
+Note that we additionally define $p_{\text{z|x}}$ as the ground-truth user feedback probability and $p_{\text{x|z}}$ as the posterior distribution of retriever scores conditioned on user feedback.
+Both $p_{\text{z|x}}$ and $p_{\text{x|z}}$ are not related to the reranker $q_{\theta}$, and they are merely determined by the retriever prior and the user.
+Then, we can extract the user behavior alignment error of $q_{\theta}$ as follows and derive its relation to the goal of $\max_{\theta}\log p(\mathbf{z}_{u})$:
 $-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u})]$ $=-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}\Big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x}_{u})}\Big]-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}[\log p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x}_{u})]$ (10)
@@
 $\Rightarrow-\log p(\mathbf{z}_{u})=$ $-\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}[\log q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u})]+\mathbb{E}_{\mathbf{x}_{u}\sim p_{\text{x}}}\Big[\log\frac{q_{\theta}(\mathbf{z}_{u}|\mathbf{x}_{u})}{p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x}_{u})}\Big]$
-$-D_{\text{KL}}\big(p_{x}(\mathbf{x}_{u})\|p_{\text{x|z}}(\mathbf{x}_{u}|\mathbf{z}_{u})\big).$
+$-D_{\text{KL}}\big(p_{\text{x}}(\mathbf{x}_{u})\|p_{\text{x|z}}(\mathbf{x}_{u}|\mathbf{z}_{u})\big).$
 which corresponds to Eq.(3).
-As we have illustrated in section 3.2, the direct optimization of the first term $\mathcal{L}_{\text{direct}}$ might be misaligned with the data likelihood maximization goal when there exists irregular retriever scores or the posterior is inconsistent with the prior distribution.
-In contrast, our proposed solution solves this limitation by collaboratively learning the noise-injecting posterior $p_{\phi}$ along with the reranker.
-The resulting objectives in Eq.(5) intentionally finds a posterior that align with the ground truth prior and explore adversarial cases that the reranker fail to accurately reconstruct the user feedback (details in section 3.4).
-Appendix B Learning Algorithm
+As we have discussed in section 3.2, the direct optimization of the first term $\mathcal{L}_{\text{direct}}$ might cause misalignment with the data likelihood maximization goal, since the observed data does not provide sufficient information about the retriever score distribution and the reranker is unaware of the noise patterns in the retriever score space.
+In contrast, our proposed solution solves this limitation by collaboratively learning the noise generator $f_{\phi}$ along with the reranker.
+The resulting objectives in Eq.(5) intentionally find a synthetic noisy score distribution $p_{\phi}$ that aligns with the real retriever score distribution and explores adversarial cases in which the reranker fails to accurately denoise into user feedback (details in Section 3.4).
+Appendix C Learning Algorithm
 We present the overall learning process of our method in Algorithm 1.
-0: Hyperparameters $\lambda_{c}$ (or $\lambda_{e}$ if using learnable posterior), $\lambda_{m}$
-0: Optimized denoising model $q_{\theta}$ (and the posterior $p_{\phi}$ if using learnable posterior).
+0: Hyperparameters $\lambda_{c}$, $\lambda_{e}$, and $\lambda_{m}$, initial reranker $\theta$ and noise generator $\phi$
+0: Optimized denoising reranker $q_{\theta}$ and the noise generator $f_{\phi}$.
 1:for Epoch $i$ do
-2:  if Using learnable posterior and $i>\lambda_{e}$ then
+2:  if $i>\lambda_{e}$ then
 3:   for Each sample $(\mathbf{u},\mathbf{x}_{u},\mathbf{z}_{u})$ do
-4:    Sample $\mathbf{x}\sim p_{\phi}^{\text{model}}$ for given $\mathbf{u}$ and $\mathbf{z}_{u}$;
-5:    Optimize $\theta$ based on $\mathcal{L}_{\theta}$;
-6:    Optimize $\phi$ based on $\mathcal{L}_{\text{adv}}$ and $\mathcal{L}_{\text{x}}^{\prime}$;
+4:    Generate $\bm{\epsilon}\sim f_{\phi}^{\text{model}}$ for given $\mathbf{u}$ and $\mathbf{z}_{u}$ and form synthetic noisy scores $\mathbf{x}_{u}^{\prime}$ with Eq (4)
+5:    Optimize reranker $\theta$ based on $\mathcal{L}_{\theta}$;
+6:    Optimize noise generator $\phi$ based on $\mathcal{L}_{\text{adv}}$ and $\mathcal{L}_{\text{x}}$;
 7:   end for
@@
 9:   for Each sample $(\mathbf{u},\mathbf{x}_{u},\mathbf{z}_{u})$ do
-10:    Sample $\mathbf{x}\sim p_{\phi}^{\text{heuristic}}$ for given $\mathbf{u}$ and $\mathbf{z}_{u}$
-11:   end for
-12:  end if
-13:end for
+10:    Sample $\bm{\epsilon}\sim f_{\phi}^{\text{heuristic}}$ for given $\mathbf{u}$ and $\mathbf{z}_{u}$ and form synthetic noisy scores $\mathbf{x}_{u}^{\prime}$ with Eq (4)
+11:    Optimize reranker $\theta$ based on $\mathcal{L}_{\theta}$;
+12:   end for
+13:  end if
+14:end for
 Algorithm 1 Training procedures of DNR
-Appendix C Detailed Settings for Offline Experiments
-C.1 Datasets
+Appendix D Detailed Settings for Offline Experiments
+D.1 Datasets
 The ML-1M dataset is a public benchmark in the field of recommender systems, comprising 1 million ratings from 6,040 users across more than 3,900 movies.
@@
 For all datasets, we initially process by eliminating users and items with fewer than 20 interactions, to avoid the cold-start issue.
-For the retriever stage, we split the data into train and test subsets at a ratio of 8:2.
+For the retriever stage, we split the data into train and test subsets at a ratio of 8:2, where each sample is composed of interaction history and an item for prediction.
 For the reranker stage, we sort the interactions in chronological order, using the last six interactions as the item list revealed to users after reranking.
@@
 Amazon-Toys 35,736 38,121 311,386
-C.2 Baselines
+D.2 Hyperparameters Setting
+Here we present the optimal hyperparameters’ settings in Table 7.
+Table 7: Performance of the retriever stage, which adopts the matrix factorization (MF) model to rate and select the top-50 items from full item set.
+Datasets ML-1M Kuaivideo Book
+$\lambda_{c}$ 0.4 0.3 0.5
+$\lambda_{m}$ 0.4 0.6 0.4
+$\lambda_{e}$ 40 80 100
+weight_decay 0.001 0.001 0
+dropout 0.3 0.3 0.5
+batch_size 2048 2048 2048
+emb_dim 128 128 128
+D.3 Baselines
 We detail the compared baselines of our main experiments in the following, including traditional recommenders, list-refinement methods, generator-evaluator methods, and diffusion-based methods.
 Traditional Recommenders: predict the scores of candidate items and rank them accordingly.
-  • SASRec [23] proposes a self-attention based sequential recommendation model (SASRec) that captures long-term semantics and short-term dynamics using an attention mechanism.
-  • Caser [36] introduces a convolutional sequence embedding recommendation model that captures sequential patterns using convolutional filters applied to latent space embeddings.
-  • GRU4Rec [37] apply recurrent neural networks (RNNs) to session-based recommendations, addressing the challenge of short user sessions without long-term profiles.
-  • MiDNN [24] proposes a global optimization framework for e-commerce search ranking that considers mutual influences between items, using extended features and sequence generation to optimize ranking and maximize overall utility.
+  • SASRec Kang & McAuley (2018) proposes a self-attention based sequential recommendation model (SASRec) that captures long-term semantics and short-term dynamics using an attention mechanism.
+  • Caser Tang & Wang (2018) introduces a convolutional sequence embedding recommendation model that captures sequential patterns using convolutional filters applied to latent space embeddings.
+  • GRU4Rec Hidasi et al. (2016) apply recurrent neural networks (RNNs) to session-based recommendations, addressing the challenge of short user sessions without long-term profiles.
+  • MiDNN Zhuang et al.
+(2018) proposes a global optimization framework for e-commerce search ranking that considers mutual influences between items, using extended features and sequence generation to optimize ranking and maximize overall utility.
 List-Refinement Methods: encode the item lists from the previous stage and refine the rankings based on this information.
-  • DLCM [8] leverages local ranking context from top-retrieved documents to refine initial ranking lists.
+  • DLCM Ai et al.
+(2018) leverages local ranking context from top-retrieved documents to refine initial ranking lists.
 The model uses a recurrent neural network to capture document interactions and an attention-based loss function, significantly improving performance over traditional learning-to-rank methods.
-  • SetRank [38] introduces a neural learning-to-rank model that learns a permutation-invariant ranking function directly on document sets.
+  • SetRank Pang et al.
+(2020) introduces a neural learning-to-rank model that learns a permutation-invariant ranking function directly on document sets.
 Utilizing multi-head self-attention blocks, SetRank captures cross-document interactions and achieves robust performance across varying input sizes.
-  • PRM [6] addresses personalized re-ranking in recommendation systems by incorporating user-specific preferences into the re-ranking process.
+  • PRM Pei et al.
+(2019) addresses personalized re-ranking in recommendation systems by incorporating user-specific preferences into the re-ranking process.
 It considers user behavior and candidate list, enhancing both relevance and personalization in the final recommendations.
-  • MIR [25] proposes to capture complex interactions between user actions and candidate list features across multiple levels, improving the accuracy and relevance of recommendations through a hierarchical reranking approach.
+  • MIR Xi et al.
+(2022) proposes to capture complex interactions between user actions and candidate list features across multiple levels, improving the accuracy and relevance of recommendations through a hierarchical reranking approach.
 Generator-Evaluator Methods: generate multiple ranked item lists and evaluate them to select the best one for users.
-  • EGRerank [22] proposes an evaluator-generator framework for learning-to-rank (LTR) in e-commerce applications.
+  • EGRerank Huzhang et al.
+(2023) proposes an evaluator-generator framework for learning-to-rank (LTR) in e-commerce applications.
 The framework includes an evaluator that evaluates recommendations involving context, a generator that maximizes the evaluator score through reinforcement learning, and a discriminator that ensures the generalization of the evaluator.
-  • Pier [9] follows a two-stage architecture with a Fine-grained Permutation Selection Module (FPSM) and an Omnidirectional Context-aware Prediction Module (OCPM).
+  • Pier Shi et al.
+(2023) follows a two-stage architecture with a Fine-grained Permutation Selection Module (FPSM) and an Omnidirectional Context-aware Prediction Module (OCPM).
 FPSM selects top-K candidate permutations based on user interest using SimHash, while OCPM evaluates these permutations with an omnidirectional attention mechanism.
-  • NAR4Rec [10] explores the use of non-autoregressive generative models for re-ranking in recommendation systems, addressing challenges such as sparse training samples and dynamic candidates by introducing a matching model, unlikelihood training to distinguish feasible sequences, and contrastive decoding to capture item dependencies.
+  • NAR4Rec Ren et al.
+(2024) explores the use of non-autoregressive generative models for re-ranking in recommendation systems, addressing challenges such as sparse training samples and dynamic candidates by introducing a matching model, unlikelihood training to distinguish feasible sequences, and contrastive decoding to capture item dependencies.
 Diffusion-Based Methods: utilize diffusion models to generate items or lists for recommendation.
-  • DiffuRec [26] incorporates a diffusion process to inject noise into target item embeddings and a reverse process to reconstruct the target item representation.
-  • DCDR [11] introduces a new framework that leverages diffusion models for the reranking stage in recommendation systems.
+  • DiffuRec Li et al. (2024) incorporates a diffusion process to inject noise into target item embeddings and a reverse process to reconstruct the target item representation.
+  • DCDR Lin et al.
+(2024) introduces a new framework that leverages diffusion models for the reranking stage in recommendation systems.
 DCDR extends traditional diffusion models with a discrete forward process and a conditional reverse process to generate high-quality item sequences.
-C.3 Implementation of Diverse Strategies for Leveraging Retriever Scores
-Here, we present the detailed implementation of different ways to leverage retriever scores in Figure C.3.
+D.4 Implementation of Diverse Strategies for Leveraging Retriever Scores
+Here, we present the detailed implementation of different ways to leverage retriever scores in Figure D.4.
 The left one refers to “c score”, which leverages a concatenate operation to combine the score features with encoded user states from users’ interaction history.
 The moderate one represents “+ score”, which incorporates the score feature with user states by a simple plus operation.
-The right one is the “w score”, which adopts the initial candidate score as the weight of candidate items.s
+The right one is the “w score”, which adopts the initial candidate score as the weight of candidate items.
 Figure 4: Different ways to leverage retriever scores.
-Appendix D Complementary Experimental Analysis
-D.1 Experiments of Retriever Stage
+Appendix E Complementary Experimental Analysis
+E.1 Experiments of Retriever Stage
 Detailed Experimental Settings for Retriever Stage
-We implement the retriever stage with the MF method [39], which aims to improve the click rate performance.
+We implement the retriever stage with the MF method Koren et al.
+(2009), which aims to improve the click rate performance.
 The dimension of the item embedding is 128, and the learning rate is 0.001.
@@
 AUC 79.20% 89.62% 88.82%
-D.2 Visualization of Learned Noises
-To visualize the injected noise distribution, we compare the density curve of the injected noise from different noise injection strategies, including Gaussian noise, Beta noise, and learning-based noise.
+E.2 Visualization of Learned Noises
+To visualize the generated noise distribution, we compare the density curve of the generated noise from different noise generators, including Gaussian noise, Beta noise, and learning-based noise.
 As demonstrated in Figure 5, the distribution of learned noise is more aligned with the distribution of true noise between retriever scores and user feedback.
 This means that the resulting posterior with model-based noise sampler better reduces the divergence term $\delta_{\text{x}}$ in Eq.(5), and potentially improve the alignment between our proposed solution with the three objectives and the data log-likelihood maximization goal.
-Figure 5: The visualization of injected noise from different variants.
-D.3 Detailed Results of Sensitivity Analysis
+Figure 5: The visualization of generaed noise from different variants.
+E.3 Performance on Different settings of Beta-Distribution
+The U-shaped Beta distribution in our paper is motivated by the binary nature of the user feedback scores and the uncertain initial guess, and it is already empirically outperforming the Gaussian distribution in our experiments.
+Since the heuristically designed beta distribution is not guaranteed to be optimal — unlike the model-based noise generator, whose learned noise distribution is demonstrated to be closer to the real noise distribution.
+To explore different Beta distributions, we conducted additional experiments with DNR that used a beta distribution Beta(2, 5), i.e., DNR-BL, which is more aligned with the true noise distribution as well as our model-based generator (as shown in Figure 5), before switching to the model-based generator in epoch $\lambda_{e}$.
+Empirically, the results are presented in Table 9.
+Table 9: Experiments with the beta distribution replicate the learned distribution.
+Methods ML-1M Kuaivideo Book
+NDCG@6 MAP@6 NDCG@6 MAP@6 NDCG@6 MAP@6
+DNR-B 77.12 67.30 70.15 58.12 83.53 75.54
+DNR-BL 77.43 67.49 71.10 59.61 83.68 75.79
+We can observe that using this more "aligned" beta distribution (i.e., , DNR-BL) before epoch $\lambda_{e}$ offers a slight improvement compared to using the U-shaped Beta (i.e., DNR-B) before $\lambda_{e}$.
+This reinforces our main claim: adopting a model-based noise generator to learn and regularize noise distribution, aligning it more closely with the ground-truth noise distribution.
+E.4 Detailed Results of Sensitivity Analysis
 Here, we present the detailed results of the sensitivity analysis for different datasets in Figure 6, Figure 7, and Figure 8.
-Figure 6: Sensitivity of DNR to the hyperparameter $\lambda_{c}$ on different datasets, which controls the degree of noise injection.
-Figure 7: Sensitivity of DNR to the hyperparameter $\lambda_{e}$ on different datasets, which decides the epochs where the adversarial learning starts.
-Figure 8: Sensitivity of DNR to the hyperparameter $\lambda_{m}$ on different datasets, which controls the magnitude of the augmentation.
-D.4 Online A/B Test
+E.5 Experiments on Larger Reranker Backbones
+DNR is a backbone-agnostic framework for reranking tasks, utilizing and denoising the noisy scores from the previous retrieval stage through adversarial learning.
+Recently, LLMs have become a powerful tool for reranking tasks.
+However, text-based LLM4Rerank is not comparable in our setting offline or online, due to the lack of rich textual information (such as detailed item descriptions) that LLM-based rerankers fundamentally rely on.
+To evaluate the effectiveness of DNR in large models, we scaled up our PRM baseline by significantly increasing the number of its Transformer layers, serving as a “larger model” for comparison.
+We then tested our DNR framework, using this scaled-up PRM as its reranker backbone.
+The results are presented in the table 10.
+Table 10: Experiments on larger models (expand transformers of PRM to 10 layers) with ML-1M dataset.
+Methods ML-1M
+HR@6 NDCG@6 MAP@6 F1@6 AUC
+original PRM 60.09 72.85 62.21 65.51 88.20
+original PRM+DNR 64.89 77.67 68.00 70.71 92.75
+larger PRM 62.95 76.09 66.72 68.61 88.68
+larger PRM+DNR 65.87 78.72 69.33 71.95 89.19
+We can observe that scaling up the backbone may improve the reranking result, comparing the larger PRM with the original PRM, or comparing the larger DNR with the original DNR.
+Additionally, our original DNR outperforms the scaled-up “larger model” baseline (larger PRM); The DNR framework may also generalize well to a larger model.
+E.6 Online A/B Test
 To better investigate the effectiveness of our DNR in an online environment, we conduct an A/B test on an industrial platform with video feeds to 100+ millions of users daily.
@@
 During the ranking process, a scoring model computes multiple dimensional features for each item.
-Among these, we utilize PCTR(predicted probability that the video will be watched for more than 3 seconds) as the primary retrieval score $\mathbf{x}_{u}$.
+Among these, we utilize PCTR(predicted probability that the video will be watched for more than 3 seconds) as the primary retrieval score $\mathbf{x}_{u}$, which corresponds to the “realshow” metric in the online system.
+Note that the final reranking stage considers the combination of the retrieval stage and the middle ranking stage as a holistic previous stage, so the output PCTR scores from the ranking stage are considered as “retrieval score” in this final stage.
 DNR is applied to generate a new set of scores for 50 selected items and the top-6 items are chosen as the final exposure.
@@
 These results indicate that the DNR method can estimate PCTR more accurately, which is strongly correlated with realshow.
-Appendix E Limitation and Discussion
-E.1 Limitations
+Additionally, the industrial recommendation environment involves various other metrics in addition to the realshow metric, including but not limited to app usage time, watch time, share-rate, like-rate, and comment-rate.
+We provide the corresponding results in Table 11.
+We can see that, in our scenario, watch-time and share-rate are slightly negatively impacted, while like-rate and comment-rate are slightly positively impacted, indicating a potential trade-off between metrics.
+In general, the impact on all other metrics is not statistically significant, indicating the effectiveness of our method.
+metric performance boost statistically significant ($p<0.05$)
+overall app-time -0.011% no
+realshow +1.089% yes
+watch-time -0.091% no
+share-rate -0.758% no
+like-rate +0.609% no
+comment-rate +0.574% no
+save-rate -0.690% no
+profile-click-rate +0.129% no
+Table 11: Extended online A/B test results.
+Figure 6: Sensitivity of DNR to the hyperparameter $\lambda_{c}$ on different datasets, which controls the degree of noise injection.
+Figure 7: Sensitivity of DNR to the hyperparameter $\lambda_{e}$ on different datasets, which decides the epochs where the adversarial learning starts.
+Figure 8: Sensitivity of DNR to the hyperparameter $\lambda_{m}$ on different datasets, which controls the magnitude of the augmentation.
+E.7 Computational Analysis
+Theoretically, our method does not change the computational cost during inference since we can adopt the same reranking model structure (e.g., PRM Pei et al.
+(2019)).
+In the training phase, we include an additional noise generator that may require O(1) computational overhead, whose cost depends on the generator’s structure.
+Empirically, the generator does not have to be sophisticated, and the computational overhead can be trivial compared to the reranker’s training.
+To better demonstrate computational cost, we compare the per-epoch training time of DNR against several baseline rerankers, including PRM Pei et al.
+(2019), Pier Shi et al.
+(2023), DCDR Lin et al.
+(2024), and a "w/ score" variant—where retriever scores are directly used as input to the reranker.
+In table 12, we present the time cost comparison for different reranker baselines in each training epoch.
+The variant “w/ score” denotes methods incorporating retriever scores as input features, regardless of the specific integration method (e.g., concatenation, plus, or weighted combination).
+Our proposed DNR demonstrates competitive training efficiency, with time costs comparable to both PRM and w/ score baselines across all datasets (ML-1M, KuaiVideo, and Book).
+This validates the computational efficiency of our approach while maintaining competitive performance.
+Table 12: Comparison of time costs.
+Datasets ML-1M Kuaivideo Book
+PRM 01m01s 02m21s 01m42s
+Pier 01m54s 04m11s 02m36s
+DCDR 02m27s 04m58s 03m02s
+w/ score 01m03s 02m49s 01m46s
+DNR 01m06s 03m12s 01m51s
+Appendix F Limitation and Discussion
+F.1 Limitations
 Computational Cost: Compared to a direct optimization framework with Eq.(2) that only learns $\theta$, our framework requires extra efforts to train a posterior model $\phi$ (if using learnable posterior) with three additional loss terms to back-propagate.
@@
 Fortunately, $\phi$ is not involved in inference, so the inference cost stays the same.
-Finding Correct Distribution Models: In this paper, we present an implementation based on the Beta distribution assumption of the prior and the posterior.
-As we have mentioned in section 3.3, finding the optimal setting for the heuristic posterior might require expert experience, and we are skeptical on whether there exists a fixed setting for all user requests.
-E.2 Future Directions
-Multi-task Scenarios: When user can engage with an item with multiple behaviors (e.g., like and share a news story), DNR may adapt to this scenario by simply formulating a multi-dimensional response space for $\mathbf{z}_{u}$ and $\mathbf{x}_{u}$.
+F.2 Future Directions
+Multi-task Scenarios: When a user can engage with an item with multiple behaviors (e.g., like and share a news story), DNR may adapt to this scenario by simply formulating a multi-dimensional response space for $\mathbf{z}_{u}$ and $\mathbf{x}_{u}$.
 Yet, different behavior signals may typically follow a heterogeneous distribution where a uniform modeling technique might become sub-optimal.
-Thus, we consider it beneficial to further investigate how the proposed framework may accommodate the multi-task solutions.
-Better Approximation for KL Divergence: In our solution, the noise regularization term $\mathcal{L}_{\text{x}}$ requires the estimation of the distribution $p_{\phi}$ and $p_{\text{x}}$.
-We approximate this by directly aligning the output scores with $\mathcal{L}_{\text{x}}$ in Eq.(8) in the implementation to avoid estimation of the prior.
-To this end, there might exist better solutions to approach the original $\mathcal{L}_{\text{x}}$ term.
-More Investigation on Beta Posterior: We have illustrated that the ground truth posterior and the prior of retriever scores are likely to be a Beta distribution.
-Yet, there still lack analytical results on the optimal personalized parameters of $\alpha$ and $\beta$ that are aligned with the ground truth.
-Further study on these parameters may reveal user behavioral patterns in the real world, which could further help improve the recommendation performance.
+Additionally, the problem may become even more sophisticated when there exist multiple retrievers that estimate the same metric, especially in industrial settings.
+In this case, one may have to come up with a more articulated integration method to clarify the connections between the user feedback and different retriever scores.
+Thus, we believe it is worth further investigating how the proposed framework may accommodate the multi-task solutions.
+Multi-stage vs.
+End-to-End: Recent advances in end-to-end recommenders have witnessed significant advances, and evidence Deng et al.
+(2025) has shown that end-to-end recommenders are potentially more effective since they no longer need to worry about the consistency challenge in the multi-stage framework.
+Yet, there is still no evidence of a general superiority comparing these two paradigms, and there is still no optimal solution for a fully consistent multi-stage recommender.
+Appendix G Use of LLMs
+Large language models (LLMs) were used only to aid writing polish—including refining sentence phrasing, logical flow, and prose clarity—without altering original meanings or technical details.
+LLMs did not participate in core research tasks (e.g., experiment design, data processing, model training, result analysis, or drafting key technical content).

### v3

@@
 We conduct extensive experiments on three public datasets and an industrial recommender system, together with analytical support, to validate the effectiveness of the proposed DNR.
+The code is released at https://github.com/maowenyu-11/DNR.
 1 Introduction
-Recommender systems aim to personalize a list of items that maximizes user engagement signals, such as purchases in e-commerce, video plays in video platforms, and link sharing in social networks.
-To achieve a feasible solution with large-scale candidate pools and user requests, industrial solutions Covington et al.
-(2016); Ricci et al.
+Recommender systems aim to generate personalized item lists and maximize users’ engagement.
+For large-scale item pools, industrial solutions Ricci et al.
 (2022); Liu et al.
-(2025) typically employ a multi-stage architecture comprising a retriever and a reranker.
-The retriever, which may further decompose into several more sub-stages of retrieval and ranking modules, efficiently filters a large item pool down to a manageable candidate set of relevant items.
-Then the reranker adopts a more sophisticated but more accurate model that refines the candidate item set into an optimal ordering for the list-wise exposure.
-To optimize the two-stage framework towards a consistent goal of recommendation (e.g., aligning true user behavior or preference) across all stages, the majority of efforts have been focusing on learning a reranker-aware retriever Gallagher et al.
+(2025) typically use a two-stage retriever-reranker architecture.
+The retriever efficiently narrows down a large item pool to a manageable relevant candidate set.
+The reranker then uses a more sophisticated model to reorder the candidates into an optimal list-wise order.
+To optimize the two-stage framework towards a consistent goal of recommendation (e.g., aligning with user behavior or preference) across all stages, the majority of efforts have been focusing on learning a reranker-aware retriever Gallagher et al.
 (2019); Xu et al.
@@
 However, there is limited work on how to achieve a retriever-aware reranker.
-Formally, the goal of the latter reranking task Carbonell & Goldstein (1998); Jin et al.
+For the latter reranker Carbonell and Goldstein (1998); Jin et al.
 (2008); Pei et al.
 (2019); Liu et al.
-(2022) is to learn an ordering function over the retrieved candidates that maximizes user responses on the top-reranked items.
-Early studies propose to re-score the initial retrieval items Jin et al.
+(2022), early studies propose to re-score the retrieval items Jin et al.
 (2008); Ai et al.
 (2018); Pei et al.
-(2019) by modeling the mutual influences among items in the candidate list.
+(2019) by modeling the mutual influences among candidate items.
 Recent state-of-the-art approaches find it more effective to formulate the reranking task as a list generation problem Shi et al.
@@
 (2024), where the lists with better list-wise rewards are given higher generation probability.
-Yet, existing methods largely overlook the potential of initial retriever scores in the input space, which may offer rich prior information to enhance the reranking performance.
-In practice, a straightforward approach to leverage initial retriever scores is to include them as input features directly, which shows promise to improve reranking performance as in Figure 1(a).
+Yet, existing methods largely overlook the potential of initial retriever scores, which may offer rich prior information for the reranker stage.
+In practice, a straightforward approach to leverage initial retriever scores is to directly include them as extra input features of the model, which shows promise to improve reranking performance as in Figure 1(a).
 Still, as we will illustrate in section 3.2, this naive solution might be far from exploiting this retriever information effectively to align with user feedback, which is addressed by the following contributions in this work.
@@
 What follows is an empirical assumption that there is higher noise in the retriever scores than the reranker scores, which naturally indicates a noise reduction process of the reranking stage.
-This noise reduction pattern is more evident when we observe the changes in the embedding representations of ground truth target items.
+This noise reduction pattern is also evident when we observe the changes in the embedding representations of ground truth target items.
 We show evidence by comparing Figure 1(c) and Figure 1(d), where the same set of retrieved candidates is not distinguished well by the retriever, but the target items are better predicted by the reranker, in terms of the embedding space.
-This further proves that the reranker behaves like a denoising model on the retriever’s noise.
-However, directly aligning the user behavior on the condition of the observed retriever score, ignoring the uncertainty of this noisy signal, may cause a misalignment with the user behavior alignment goal in terms of the entire system.
+The reranker’s noise is more concentrated near zero than the retriever’s (Figure 1(e)), further confirming the reranker acts as a denoising model for the retriever’s noise.
+However, due to the uncertainty inherent in noisy retriever scores, denoising based on these scores may misalign with the system-wide goal of aligning with user behavior.
 Thus, we argue that learning a noise-aware reranker is crucial for achieving a retriever-aware reranker.
@@
 (b) Performance comparison of two stages in MovieLens dataset.
-(c) Target item and selected candidates embedding in the retriever stage.
-(d) Noise reduction process in the item embedding perspective.
+(c) Target item and selected candidates embedding in retriever.
+(d) Noise reduction process in item embedding perspective.
+(e) Noise density distribution of the reranker and retriever.
 Figure 1: (a) The informative retriever score and (b-d) noise reduction nature of reranking.
-The reranker in this example uses transformer model to select top-20 items from the candidates.
-“Rerank w/ s” represents using retriever scores as additional item features for reranker.
-The blue shaded circles in (c) and (d) represent the retriever’s selection of candidate items.
-The green shaded circle in (d) represents the reranker’s selection of the items of exposure.
+The reranker in this example uses transformer model to select the top-20 items from the candidates.
+“Rerank w/ s” represents using retriever scores as additional item features for the reranker.
+The black shaded circles in (c) and (d) represent the retriever’s selection of candidate items, and the green shaded circle in (d) represents those of the reranker for exposure.
+(e) compares the noise distribution (distance between predicted scores and ground-truth labels) of the retriever and the reranker.
 The proposed solution: Motivated by the previous analysis, we propose an adversarial framework, which formulates a denoising neural reranker (DNR) with a carefully designed noise generation module, which introduces unseen noise to users’ feedback to augment the distribution modeling of retriever scores.
@@
 While most existing work focuses on independently optimizing individual stages (e.g., retrieval Wang et al.
-(2011; 2017) or reranking stage Liu et al.
+(2011; 2017) or the reranking stage Liu et al.
 (2022); Pei et al.
@@
 (2024a).
-However, our work takes a complementary direction—–using earlier-stage scores to regulate later-stage optimization.
+In contrast, our work takes a complementary direction—–using earlier-stage scores to regulate later-stage optimization.
 2.2 Reranking Strategies
@@
 Existing re-ranking approaches can be broadly categorized into the following paradigms.
-Single-point re-scoring methods Kang & McAuley (2018); Zhuang et al.
+Single-point re-scoring methods Kang and McAuley (2018); Zhuang et al.
 (2018) independently predict a refined score for each item and re-rank them accordingly.
@@
 Techniques such as Transformer-based architectures Pei et al.
-(2019) capture user preference to optimize the list holistically.
+(2019) capture user preferences to optimize the list holistically.
 Generator-evaluator methods Huzhang et al.
@@
 (2023) first generate multiple candidate lists and then evaluate them using a learned utility function to select the best one.
-They can effective generate a high-quality recommendation list but suffer from the increasing computational cost due to the generation-evaluation loop.
-Emerging techniques leverage diffusion models Lin et al.
-(2024); Li et al.
-(2024) to generate refined item lists by iteratively denoising a ranking distribution, which excel in capturing complex user preferences.
+They can effectively generate a high-quality recommendation list but suffer from the increasing computational cost due to the generation-evaluation loop.
+Emerging techniques leverage diffusion models Mao et al.
+(2025c; a; d) to generate refined item lists by iteratively denoising a ranking distribution, which excels in capturing complex user preferences.
 2.3 Adversarial Learning
-Adversarial learning Lowd & Meek (2005) has emerged as a powerful technique for data augmentation in recommendation systems, improving the models robustness and generalization.
+Adversarial learning Lowd and Meek (2005) has emerged as a powerful technique for data augmentation in recommendation systems, improving the model robustness and generalization.
 By leveraging generative adversarial networks (GANs) Wang et al.
@@
 For instance, IRGAN Wang et al.
-(2017) employs a minimax game between a generator (sampling hard negative items) and a discriminator (distinguishing real vs.
-synthetic interactions) to improve ranking performance.
-Similarly, AdvIR Park & Chang (2019) introduces adversarial perturbations to embedding spaces, forcing the model to learn more generalizable representations.
+(2017) employs a minimax game between a generator (sampling hard negative items) and a discriminator (distinguishing real from synthetic interactions) to improve ranking performance.
+Similarly, AdvIR Park and Chang (2019) introduces adversarial perturbations to embedding spaces, forcing the model to learn more generalizable representations.
 These methods demonstrate that adversarial learning can not only augment training data but also enhance models’ robustness.
-Our method is designed to add noise by adversarial learning, which can enhance the denoising models (rerankers)’ effectiveness in denosing the pre-ranking scores to align with user feedback.
+Our method is designed to add noise by adversarial learning, which can enhance the reranker’s effectiveness in denoising the pre-ranking scores to align with user feedback.
 3 Methodology
@@
 $L_{2}$ $=-D_{\text{KL}}\big(p_{\text{x}}(\mathbf{x})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big).$
-where $p_{\text{z|x}}$ represents the ground-truth user feedback probability and $p_{\text{x|z}}$ represents the posterior distribution of retriever scores conditioned on user feedback, both terms are not related to the reranker $q_{\theta}$ and they are merely determined by the retriever prior and the user.
+where $p_{\text{z|x}}$ represents the ground-truth user feedback probability and $p_{\text{x|z}}$ represents the posterior distribution of retriever scores conditioned on user feedback, both terms are not related to the reranker $q_{\theta}$, and they are merely determined by the retriever prior and the user.
 Intuitively, smaller values of $L_{1}$ and $L_{2}$ make the solution of Eq.(2) more aligned with the goal of Eq.(1).
@@
 To overcome the aforementioned limitations, we propose to complement the denoising reranker $q_{\theta}$ with a noise generator $f_{\phi}(\cdot|\mathbf{z}_{u})$.
-Formally, we generate noise with $\bm{\epsilon}\sim f_{\phi}$ and add it to the user feedback to form the synthetic retriever score posterior $p_{\phi}(\mathbf{x}|\mathbf{z}_{u})$, taking insight from the reparameterization trick Kingma & Welling (2014):
+Formally, we generate noise with $\bm{\epsilon}\sim f_{\phi}$ and add it to the user feedback to form the synthetic retriever score posterior $p_{\phi}(\mathbf{x}|\mathbf{z}_{u})$, taking insight from the reparameterization trick Kingma and Welling (2014):
 $\mathbf{x}_{u}^{\prime}=(1-\lambda_{c})\mathbf{z}_{u}+\lambda_{c}\bm{\epsilon},$ (4)
@@
 (2024), but we notice that this is not a good representation of the noise in retriever scores.
-Intuitively, given the binary user feedback signals, a natural assumption of the retriever prior is the Beta distribution, where we can derive the heuristic solution as $\bm{\epsilon}\sim\text{Beta}(\alpha,\beta)$.
+Intuitively, given the binary user feedback signals, a natural assumption of the retriever prior is the Beta distribution, i.e., , the conjugate prior and the posterior for binary variables (i.e., bernoulli or binomial distribution) are both beta distribution, where we can derive the heuristic solution as $\bm{\epsilon}\sim\text{Beta}(\alpha,\beta)$.
 For simplicity, one may set $\alpha=\beta=0.5$ to approximate the uncertainty.
 In practice, we find that it is also beneficial to set $\alpha$ and $\beta$ that align with the observed noise distribution in the data.
-  • Model-based Generator: While the heuristic generator is effective in our empirical study, we are skeptical of whether the heuristic noise generator is optimal for augmenting the retriever score distribution.
+  • Model-based Generator: While the heuristic generator is effective in our empirical study, we are skeptical of whether the manually designed heuristic noise generator is optimal for augmenting the retriever score distribution.
+Since the prior noise distribution is uncertain, it is necessary to learn this distribution adaptively and personally.
 To achieve a more precise estimation, we design a learnable model $f_{\phi}^{\text{model}}$, with parameter $\phi$, that will be optimized along with the reranker (details in the next section).
 The generated noise $\bm{\epsilon}$ should also be aligned with the observed distribution, and it also participates in an adversarial training objective in order to better explore the retriever score space, which in turn improves the reranker’s effectiveness.
+As the noise is learned, $\lambda_{c}$ requires no manual adjustment and can be treated as 1.
 As we will show in Section 4.3, this choice is empirically better than the heuristic generator.
 3.4 Overall Learning Framework
-In this section, we show how to optimize the user feedback alignment goal i.e., Eq.(1) by simultaneously learning the denoising reranker and the adversarial noise generator.
+In this section, we show how to optimize the user feedback alignment goal, i.e., Eq.(1) by simultaneously learning the denoising reranker and the adversarial noise generator.
 Specifically, we can decompose the negative log-likelihood objective into three loss terms (derivation in Appendix B.1):
@@
 We implement the model-based generator $f_{\phi}^{\text{model}}$ by a 2-layer MLP and directly output $\mathbf{x}_{u}^{\prime}$ using Eq.(4), so that the end-to-end learning becomes feasible for Eq.(7).
-In cases where the noise generator is not directly optimizable, one may have to consider other alternatives like reparameterization tricks Kingma & Welling (2014) or policy gradient Silver et al.
+In cases where the noise generator is not directly optimizable, one may have to consider other alternatives like reparameterization tricks Kingma and Welling (2014) or policy gradient Silver et al.
 (2014).
@@
 In this section, we conduct extensive experiments to answer the following questions:
-RQ1: How does DNR perform in the reranking tasks of recommendation systems compared with leading rerankers?
-RQ2: How well is the denoising formulation compared to alternative methods of leveraging retriever scores?
-RQ3: Can the denoising formulation generalize to other reranker backbones, rather than PRM?
-RQ4: What is the performance difference between the heuristic noise generator and the adversarial model-based generator?
+  • RQ1: How does DNR perform compared with leading rerankers in recommendation systems?
+  • RQ2: How does the denoising formulation compare to other methods of using retriever scores?
+  • RQ3: Can the denoising formulation generalize to reranker backbones other than PRM?
+  • RQ4: How does the heuristic noise generator’s performance compare to that of the adversarial model-based generator?
   • RQ5: How does different objectives in DNR contribute to the performance?
@@
 Datasets.
-We conduct experiments on three real-world datasets, including ML-1M Harper & Konstan (2016), Kuaivideo Liu et al.
+We conduct experiments on three real-world datasets, including ML-1M Harper and Konstan (2016), Kuaivideo Liu et al.
 (2025), and Amazon-books Ni et al.
@@
 For the online experimental verification, please refer to Appendix E.6.
-Baselines: We compare the performance of DNR with four categories of recommendation methods in our reranking task, including traditional recommenders (i.e., SASRec Kang & McAuley (2018), Caser Tang & Wang (2018), GRU4Rec Hidasi et al.
+Baselines: We compare the performance of DNR with four categories of recommendation methods in our reranking task, including traditional recommenders (i.e., SASRec Kang and McAuley (2018), Caser Tang and Wang (2018), GRU4Rec Hidasi et al.
 (2016), and MiDNN Zhuang et al.
@@
 NAR4Rec 62.81 75.01 65.42 68.31 88.30 44.31 63.83 51.45 52.08 89.94 70.08 79.46 70.69 72.66 92.44
+MG-E 63.53 76.18 66.52 69.37 89.08 46.89 66.19 53.75 55.26 90.16 70.63 79.81 71.07 73.20 90.21
 DiffuRec 62.16 75.80 66.28 67.83 91.48 40.78 61.16 50.77 48.00 81.08 58.62 68.77 57.25 60.79 84.74
@@
 We conduct sensitivity analysis for $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$ for both DNR-G and DNR-B, and present the results in Figure 3.
+When adjusting $\lambda_{m}$, $\lambda_{e}$ is fixed at 40; when adjusting $\lambda_{e}$, $\lambda_{m}$is fixed at 0.3.
+For $\lambda_{c}$ adjustments, $\lambda_{m}$ and $\lambda_{e}$ are set to 0.6 and 80, respectively.
 While the Beta noise generally performs better than Gaussian noise in most settings, as shown in Table 2, there exist some common patterns in these hyper-parameters: 1) all parameters appears to have an optimal value within the searched region, indicating the effectiveness of including each corresponding designs, i.e., $\lambda_{c}$ for the synthetic noisy scores in Eq.(4), $\lambda_{m}$ for the combined denoising loss Eq.(6), and $\lambda_{e}$ for the switching from heuristic noise generator to model-based one; 2) Over-amplifying these hyper-parameters results in the downgrade of the rerank performance, indicating the significance of keeping a balance between the involved components.
@@
 Theoretical analysis and empirical validation across three public datasets demonstrate the effectiveness of the noise reduction formulation and the design of our DNR.
+6 Acknowledgement
+This research was supported by the National Natural Science Foundation of China (U25A20445) and the advanced computing resources provided by the Supercomputing Center of the University of Science and Technology of China (USTC).
+This work was also supported by Kuaishou Technology, where author Wenyu Mao conducted this research during an internship.
 Ethics Statement.
@@
 Reproducibility Statement.
-All the results in this work are reproducible. We have discussed the optimal hyperparameters and the details on devices and software environments in Section 4.1 and Appendix D.2.
+All the results in this work are reproducible.
+We have discussed the optimal hyperparameters and the details on devices and software environments in Section 4.1 and Appendix D.2.
+For reproducibility, we release the code at https://github.com/maowenyu-11/DNR
 Appendix A Notations
@@
 In this section, we illustrate the derivation of the Eq.(5).
-Similar to the variational Bayesian inference Kingma & Welling (2014), we start by analyzing the KL divergence between the real retriever scores and synthetic noisy scores:
+Similar to the variational Bayesian inference Kingma and Welling (2014), we start by analyzing the KL divergence between the real retriever scores and synthetic noisy scores:
 $D_{\text{KL}}(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u}))$ $=-\mathbb{E}_{p_{\phi}}\Big[\log\frac{p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})}{p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}\Big]$ (8)
@@
 $=-\mathbb{E}_{p_{\phi}}\Big[\frac{\log p(\mathbf{x},\mathbf{z}_{u})}{\log p_{\phi}(\mathbf{x}|\mathbf{z}_{u})}\Big]+\log p(\mathbf{z}_{u}),$
-Note that a key difference between Eq.(8) and the conventional derivation in Kingma & Welling (2014) is that the variational posterior $\phi$ is not a latent factor encoder that reduces noise and dimensions, but an explicit noise generator that mimics the behavior of retriever scores.
+Note that a key difference between Eq.(8) and the conventional derivation in Kingma and Welling (2014) is that the variational posterior $\phi$ is not a latent factor encoder that reduces noise and dimensions, but an explicit noise generator that mimics the behavior of retriever scores.
 In terms of the denoising problem formulation, it is also similar to a one-step diffusion, but with a key difference in that the real retriever scores $\mathbf{x}_{u}$ are not sampled by a manually defined distribution.
@@
 Here we present the optimal hyperparameters’ settings in Table 7.
-Table 7: Performance of the retriever stage, which adopts the matrix factorization (MF) model to rate and select the top-50 items from full item set.
+Table 7: Hyperparameter setting for DNR on different datasets
 Datasets ML-1M Kuaivideo Book
@@
 Traditional Recommenders: predict the scores of candidate items and rank them accordingly.
-  • SASRec Kang & McAuley (2018) proposes a self-attention based sequential recommendation model (SASRec) that captures long-term semantics and short-term dynamics using an attention mechanism.
-  • Caser Tang & Wang (2018) introduces a convolutional sequence embedding recommendation model that captures sequential patterns using convolutional filters applied to latent space embeddings.
+  • SASRec Kang and McAuley (2018) proposes a self-attention based sequential recommendation model (SASRec) that captures long-term semantics and short-term dynamics using an attention mechanism.
+  • Caser Tang and Wang (2018) introduces a convolutional sequence embedding recommendation model that captures sequential patterns using convolutional filters applied to latent space embeddings.
   • GRU4Rec Hidasi et al. (2016) apply recurrent neural networks (RNNs) to session-based recommendations, addressing the challenge of short user sessions without long-term profiles.
@@
 This means that the resulting posterior with model-based noise sampler better reduces the divergence term $\delta_{\text{x}}$ in Eq.(5), and potentially improve the alignment between our proposed solution with the three objectives and the data log-likelihood maximization goal.
-Figure 5: The visualization of generaed noise from different variants.
+Figure 5: The visualization of generated noise from different variants.
 E.3 Performance on Different settings of Beta-Distribution
@@
 DNR is a backbone-agnostic framework for reranking tasks, utilizing and denoising the noisy scores from the previous retrieval stage through adversarial learning.
-Recently, LLMs have become a powerful tool for reranking tasks.
+Recently, LLMs have become a powerful tool for reranking tasks Mao et al.
+(2025b).
 However, text-based LLM4Rerank is not comparable in our setting offline or online, due to the lack of rich textual information (such as detailed item descriptions) that LLM-based rerankers fundamentally rely on.
@@
 To better investigate the effectiveness of our DNR in an online environment, we conduct an A/B test on an industrial platform with video feeds to 100+ millions of users daily.
-Detailed Experimental Settings for Online Test The multi-stage pipeline of online system consists of three stages: retrieval, ranking, and reranking.
+Detailed Experimental Settings for Online Test The multi-stage pipeline of the online system consists of three stages: retrieval, ranking, and reranking.
 In the final reranking stage, the top-6 items for user exposure are selected from a candidate set of 50 items generated by the ranking stage.
-During the ranking process, a scoring model computes multiple dimensional features for each item.
+During the ranking process, a scoring model computes multi-dimensional features for each item.
 Among these, we utilize PCTR(predicted probability that the video will be watched for more than 3 seconds) as the primary retrieval score $\mathbf{x}_{u}$, which corresponds to the “realshow” metric in the online system.
@@
 Each experiment is conducted for at least 7 days to ensure the reliability of the results.
-The DNR method demonstrates a significant improvement in realshow(realshow refers to the cumulative number of videos watched by users.), increasing it by 1.089% .
+The DNR method demonstrates a significant improvement in realshow(, i.e., the cumulative number of videos watched by users), increasing it by 1.089%.
 These results indicate that the DNR method can estimate PCTR more accurately, which is strongly correlated with realshow.
@@
 We can see that, in our scenario, watch-time and share-rate are slightly negatively impacted, while like-rate and comment-rate are slightly positively impacted, indicating a potential trade-off between metrics.
-In general, the impact on all other metrics is not statistically significant, indicating the effectiveness of our method.
+Notably, while the watch time exhibited a slight decrease, we additionally observed that the long-term DAU slightly increased by 0.01% [-0.012%,+0.012%], which is nearly reaching a statistically significant improvement.
+This potentially suggests that the slight reduction in viewing duration is likely attributable to users exploring a broader range of clicked videos, while more significant long-term indicators remain positive.
+Beyond these, the impact on all other metrics is not statistically significant, which validates the effectiveness of our method.
+It is worth noting that achieving performance gains across all online metrics is inherently a multi-objective challenge—an area that remains a key focus for future work.
+Additionally, we believe applying DNR to other xTR signals (e.g., watch time and share-rate) holds significant potential for further improvements.
 metric performance boost statistically significant ($p<0.05$)
@@
 profile-click-rate +0.129% no
+Long term DAU +0.01 % yes
 Table 11: Extended online A/B test results.
@@
 Figure 8: Sensitivity of DNR to the hyperparameter $\lambda_{m}$ on different datasets, which controls the magnitude of the augmentation.
-E.7 Computational Analysis
-Theoretically, our method does not change the computational cost during inference since we can adopt the same reranking model structure (e.g., PRM Pei et al.
-(2019)).
-In the training phase, we include an additional noise generator that may require O(1) computational overhead, whose cost depends on the generator’s structure.
-Empirically, the generator does not have to be sophisticated, and the computational overhead can be trivial compared to the reranker’s training.
-To better demonstrate computational cost, we compare the per-epoch training time of DNR against several baseline rerankers, including PRM Pei et al.
-(2019), Pier Shi et al.
-(2023), DCDR Lin et al.
-(2024), and a "w/ score" variant—where retriever scores are directly used as input to the reranker.
-In table 12, we present the time cost comparison for different reranker baselines in each training epoch.
-The variant “w/ score” denotes methods incorporating retriever scores as input features, regardless of the specific integration method (e.g., concatenation, plus, or weighted combination).
-Our proposed DNR demonstrates competitive training efficiency, with time costs comparable to both PRM and w/ score baselines across all datasets (ML-1M, KuaiVideo, and Book).
-This validates the computational efficiency of our approach while maintaining competitive performance.
-Table 12: Comparison of time costs.
-Datasets ML-1M Kuaivideo Book
-PRM 01m01s 02m21s 01m42s
-Pier 01m54s 04m11s 02m36s
-DCDR 02m27s 04m58s 03m02s
-w/ score 01m03s 02m49s 01m46s
-DNR 01m06s 03m12s 01m51s
-Appendix F Limitation and Discussion
-F.1 Limitations
-Computational Cost: Compared to a direct optimization framework with Eq.(2) that only learns $\theta$, our framework requires extra efforts to train a posterior model $\phi$ (if using learnable posterior) with three additional loss terms to back-propagate.
-This might multiply the training cost for several folds.
-Fortunately, $\phi$ is not involved in inference, so the inference cost stays the same.
-F.2 Future Directions
-Multi-task Scenarios: When a user can engage with an item with multiple behaviors (e.g., like and share a news story), DNR may adapt to this scenario by simply formulating a multi-dimensional response space for $\mathbf{z}_{u}$ and $\mathbf{x}_{u}$.
-Yet, different behavior signals may typically follow a heterogeneous distribution where a uniform modeling technique might become sub-optimal.
-Additionally, the problem may become even more sophisticated when there exist multiple retrievers that estimate the same metric, especially in industrial settings.
-In this case, one may have to come up with a more articulated integration method to clarify the connections between the user feedback and different retriever scores.
-Thus, we believe it is worth further investigating how the proposed framework may accommodate the multi-task solutions.
-Multi-stage vs.
-End-to-End: Recent advances in end-to-end recommenders have witnessed significant advances, and evidence Deng et al.
-(2025) has shown that end-to-end recommenders are potentially more effective since they no longer need to worry about the consistency challenge in the multi-stage framework.
-Yet, there is still no evidence of a general superiority comparing these two paradigms, and there is still no optimal solution for a fully consistent multi-stage recommender.
-Appendix G Use of LLMs
-Large language models (LLMs) were used only to aid writing polish—including refining sentence phrasing, logical flow, and prose clarity—without altering original meanings or technical details.
-LLMs did not participate in core research tasks (e.g., experiment design, data processing, model training, result analysis, or drafting key technical content).
+E.7 Experiments on Diverse Retrievers
+To better illustrate the generalizability of DNR across different retrievers, we implemented the dual-tower (DT) model as a stronger retriever.
+The comparison results in the retrieval stage are summarized in Table 12, and the comparison results in the reranking stage are summarized in Table 13.
+We can see that DT+DNR outperforms DT+PRM, validating the generalizability of our method and the sustained advantage of the denoising framework.
+Theoretically, there will always be an empirical error that upper bounds the overall recommender system, which aims to minimize the "optimizable error", and there will be no room for the reranker to "denoise the scores" if the retriever itself achieve this bound.
+Yet, this theoretical bound is still impractical for a single retriever in most industrial platforms, due to various reasons including (but not limited to) the computational bottleneck, multi-task trade-off, and filtering on large-scale dynamic candidate pool.
+A more advanced retriever means a smaller remaining error that can be optimized, which indicates that the reranker will solve a simpler denoising problem and simpler reranker design.
+In this case, a diminishing marginal gain only indicates a diminishing computational requirement in the reranking stage.
+In general, the advantage of the denoising formulation is maintained regardless of the retriever employed.
+Table 12: The AUC performance of the different retrievers, including the matrix factorization (MF) model and dual-tower model.
+ML-1M Kuaivideo Book
+MF 79.20% 89.62% 88.82%
+DT 79.86% 90.38% 90.0%
+Table 13: The performance comparison of rerankers on different retrievers.
+Methods ML-1M Kuaivideo Book
+NDCG@6 MAP@6 NDCG@6 MAP@6 NDCG@6 MAP@6
+MF+PRM 72.85 62.21 55.93 42.97 76.88 66.44
+MF+PIER 75.99 65.98 65.11 52.55 80.22 71.62
+MF+DNR 77.67 68.00 70.15 58.12 83.53 75.54
+DT+PRM 76.68 67.37 65.77 53.41